### v2

# Monotone Lipschitz-Gradient Denoiser: Explainability of Operator Regularization Approaches Free From Lipschitz Constant Control

Abstract
This paper addresses explainability of the operator-regularization approach under the use of monotone Lipschitz-gradient (MoL-Grad) denoiser — an operator that can be expressed as the Lipschitz continuous gradient of a differentiable convex function. We prove that an operator is a MoL-Grad denoiser if and only if it is the “single-valued” proximity operator of a weakly convex function. An extension of Moreau’s decomposition is also shown with respect to a weakly convex function and the conjugate of its convexified function. Under these arguments, two specific algorithms, the forward-backward splitting algorithm and the primal-dual splitting algorithm, are considered, both employing MoL-Grad denoisers. These algorithms generate a sequence of vectors converging weakly, under conditions, to a minimizer of a certain cost function which involves an “implicit regularizer” induced by the denoiser. Unlike the previous studies of operator regularization, our framework requires no control of the Lipschitz constant in learning the denoiser. The theoretical findings are supported by simulations.
Index Terms:
weakly convex function, proximity operator, nonexpansive operator, convex optimization
I Introduction
Operator splitting (a.k.a. proximal splitting) [1] is a standard technique nowadays for nonsmooth convex optimization. It allows to take into account prior knowledge such as sparsity and low-rankness both of which are quite relevant to signal processing as well as machine learning. Indeed, the naive discrete-valued measures (the $\ell_{0}$ pseudo-norm and the matrix rank) of such prior information are discontinuous, and one often resorts to their convex relaxations (the $\ell_{1}$ and nuclear norms) which are typically nonsmooth. Moreau’s proximity operator resides at the core of operator splitting algorithms, and it has many examples ranging from the soft-shrinkage and the convex projection to activation functions including rectified linear unit (ReLU), sigmoid, and softmax [2], to name a few. One of the remarkable advantages of the operator splitting algorithm is its computational efficiency as long as the involved nonsmooth functions are prox friendly. The celebrated examples of the operator splitting algorithm include the proximal forward-backward splitting method (the proximal gradient method) [3, 4], the Douglas-Rachford splitting method [3], the alternating direction methods of multipliers [5, 6], and the primal-dual splitting methods [7, 8, 9]. See [10, 11], for instance, for more about the history of the operator splitting algorithms.
Designing an objective function involving nonsmooth functions is the first step in typical signal processing approaches based on operator splitting. Then, an algorithm to minimize the objective is either constructed or selected from the available ones, followed by a derivation of Moreau’s proximity operators, as well as other operators such as the gradient in most cases. As the nonsmooth functions, the convex relaxations mentioned in the previous paragraph are usually chosen as a first choice. Despite the mathematical tractability, such convex regularizers tend to yield serious estimation biases [12] owing to the convexity. One might therefore return to the first step to devise a better objective function. Indeed, nonconvex regularizers have been studied widely [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], motivated by the fact that those regularizers are better approximations of the ideal $\ell_{0}$ pseudo-norm and therefore reduce the estimation bias significantly.
Some promising approaches combining convex and nonconvex functions have been studied [26, 27, 28]. In these approaches, a weakly convex penalty is coupled with a strongly convex fidelity so that convexity of the whole objective function is preserved. A generalized penalty function has been proposed in [29] so that overall convexity can be preserved when applied to inverse problems of underdetermined linear systems (where the fidelity function is not strongly convex). In [30], the penalty has further been generalized to permit a composition with a bounded linear operator, and the generalized penalty has been shown to improve the performance of the $\ell_{1}$-based total variation as a particular instance. In [31], the convex-nonconvex strategy has been applied to outlier-robust signal recovery, and a general framework has been established to clarify when the sufficient condition of overall convexity is also a necessary condition, i.e., when the bound of the regularization parameter is tight. Its extension has also been presented in [32].
There are several other related approaches. One is to directly devise a better “operator” that accommodates prior information in such an efficient way not to cause serious estimation biases (see [33] and the references therein). Another popular one is an operator-regularization approach called plug-and-play [34], where the proximity operator is replaced by another denoiser such as deep neural networks. “Explainability” of the plug-and-play method has been studied actively in recent years (see Section V-C). Most of the previous works in this line of research impose the strong assumption of (averaged) nonexpansiveness on the neural network, which often requires additional costs [35, 36]. There are some works considering a weaker assumption allowing the Lipschitz constant to exceed one [37, 38]. Whereas all those works concern convergence of the iterates, studies on optimality of the limit point are rather limited. In the present work, the following question is addressed for operators which are not necessarily nonexpansive: when do the operator-regularization approaches actually minimize some sort of objective function, and how can such a function be characterized?
In our approach, we suppose that an operator is given. It could be designed by human, or it could also be learned by machine. The designed/learned operator replaces the proximity operator appearing in an operator splitting algorithm, and the optimization problem associated with the resultant algorithm is “derived” subsequently. As such, the processes of “design” and “derivation” are in the reverse order compared to the typical optimization-based approaches. We clarify that the estimate is not explicitly characterized as a minimizer (nor a stationary/critical point) of a user-designed function. It can only be characterized as a minimizer of a convex function involving an implicit regularizer induced by a user-designed “operator”.111 It can also be characterized as a fixed point of some averaged nonexpansive operator from a wider viewpoint. A practical advantage of the proposed approach is that the operator can be designed in a way that it can be computed simply but its corresponding cost function could be complicated, or even impossible to express in a closed form.
To present a more concrete picture, let us consider the simple iterate:
$x_{k+1}:=T(x_{k}-\mu\nabla f(x_{k})),~{}k\in{\mathbb{N}},$ (1)
where $T$ is a nonlinear operator from a real Hilbert space ${\mathcal{H}}$ to ${\mathcal{H}}$, $f:{\mathcal{H}}\rightarrow{\mathbb{R}}$ is a smooth convex function, and $\mu\in{\mathbb{R}}_{++}$ is the step size. Here, the term smooth (or $\kappa$-smooth more specifically) is used in this paper to mean that the function is differentiable with its gradient Lipschitz continuous with constant $\kappa>0$ (see (3)) over ${\mathcal{H}}$. If in particular $T$ is Moreau’s proximity operator of a proper lower semicontinuous convex function $g$, (1) is the classical forward-backward splitting algorithm to minimize $\mu f+g$. It will turn out that $(x_{k})_{k\in{\mathbb{N}}}$ converges weakly to a minimizer (if exists) of a certain cost function under the following condition (in addition to other technical assumptions).222 One may consider such a condition that $T$ is a selection of the subdifferential of a (not necessarily differentiable) proper lower-semicontinuous convex function $\psi$. However, this seemingly weaker condition together with Lipschitz continuity implies Fréchet differentiability of $\psi$ [39, Proposition 17.41]. To the best of authors’ knowledge, the subdifferentiability is often used in convex analysis to relate an operator to optimization. The case of $\beta\geq 1$ corresponds to Moreau’s proximity operator of convex function [40], thus excluded in the present study.
Definition 1 (MoL-Grad Denoiser)
An operator $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ is said to be a monotone Lipschitz-gradient (MoL-Grad) denoiser if it can be expressed as the gradient of a $\beta^{-1}$-smooth convex function for $\beta\in(0,1)$; i.e., $T=\nabla\psi$ for a Fréchet differentiable convex function $\psi$ with the $\beta^{-1}$-Lipschitz continuous gradient $\nabla\psi$.
The term “denoiser” is used in Definition 1 rather than the general term “operator”, because we primarily consider denoising operator (which is often referred to as denoiser in the literature) for $T$. The term “monotone” stems from the fact that convexity of $\psi$ can be characterized by monotonicity of the gradient $\nabla\psi$ (see Section II-D). The convergence argument can be extended to other operator splitting algorithms than (1), as elaborated later on. It should be mentioned that (i) the Lipschitz continuity imposed on MoL-Grad denoiser is weaker than (averaged) nonexpansiveness which requires the Lipschitz constant to be one (or smaller), (ii) every MoL-Grad denoiser is $\beta$-cocoercivity of $T$, i.e., $\beta T$ is firmly nonexpansive, and (iii) the present study includes nonseparable operators in scope.333 For instance, when used as a penalty function for the linear inverse problem with underdetermined systems, a nonconvex function must be nonseparable to preserve overall convexity; i.e., there exists no separable nonconvex function that makes the overall cost function convex [41]. We mention that, when a given denoiser $T:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ is a continuously differentiable mapping, it can be expressed as $T=\nabla h$ for some $h:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ (i.e., $T$ is a conservative vector field) if and only if $T$ possesses a symmetric Jacobian matrix at every point [42].
I-A Contributions
In the first main part (Section III), we link the operator $T$ to a generalized notion of Moreau’s proximity operator for possibly nonconvex functions to facilitate the analysis. Here, the proximity operator of a nonconvex function is set-valued in general. In the meanwhile, every MoL-Grad denoiser is continuous owing directly to its Lipschitz continuity (see Definition 3 in Section II-C); if $T$ is discontinuous, the convergence analysis of the algorithm in (1) would hardly be tractable in general. As long as imposing continuity on $T$, we can restrict ourselves to the single-valued case, as seen from the following proposition.
Proposition 1
Given a proper function $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$, let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ be a continuous444 Continuity of the denoiser is often assumed in the literature [43]. operator such that $T(x)\in\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}[f(y)+(1/2)\left\|x-y\right\|^{2}]\neq\varnothing$ for every $x\in{\mathcal{H}}$. Then, the penalized function $f+(1/2)\left\|x-\cdot\right\|^{2}$ has a unique minimizer for every $x\in{\mathcal{H}}$.
Proof.
See Appendix A. ∎
Proposition 1 states that, if $f+(1/2)\left\|x-\cdot\right\|^{2}$ has multiple minimizers, there is no hope to find a continuous operator $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ in the form of $T(x)\in\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}[f(y)+(1/2)\left\|x-y\right\|^{2}]$. We therefore define our generalized proximity operator as a unique minimizer of the penalized function only when the minimizer exists uniquely. Now, a natural question is what is a sufficient condition for $T$ to be continuous. It is clear that $f+(1/2)\left\|x-\cdot\right\|^{2}$ has a unique minimizer (i.e., the necessary condition in the proposition holds true) if $f+(1/2)\left\|\cdot\right\|^{2}$ is strongly convex; $f$ is weakly convex at least in this case. This condition actually implies that the denoiser $T$ can be expressed as the Lipschitz continuous gradient of a convex function, and thus $T$ is continuous. Our findings of the first part are summarized below.
  1. 1.
    An operator $T$ is a MoL-Grad denoiser if and only if $T$ is the (generalized) proximity operator of a ($1-\beta$)-weakly convex function $\varphi=\psi^{*}-(1/2)\left\|\cdot\right\|^{2}$ (Theorem 1). This actually motivates the definition of MoL-Grad denoiser.
  2. 2.
    Moreau’s decomposition is extended to weakly convex functions (Propositions 2 and 3).
In the second part (Sections IVVI), we consider the modified operator splitting algorithms with the proximity operator replaced by a MoL-Grad denoiser, analyzing the convergence based on the results established in the first part. As a case study, we highlight two specific algorithms: (i) the forward-backward splitting type algorithm given in (1), and (ii) the primal-dual splitting type algorithm which is basically the Condat–Vũ algorithm [8, 9] but with a certain modification. The specific contributions are listed below.
  1. 3)
    The iterate (1) with a MoL-Grad denoiser $T$ converges weakly to a minimizer of $\mu f+\varphi$ under conditions (Theorem 2). The monotonicity of our denoiser $T$ allows its associated regularizer $\varphi$ to be (weakly) convex.
  2. 4)
    The primal-dual splitting type algorithm employing a MoL-Grad denoiser converges weakly to a minimizer of a certain function under conditions (Theorem 3).
  3. 5)
    A systematic way of building operator-regularization algorithms using MoL-Grad denoisers with convergence guarantee is presented (Section IV-C).
  4. 6)
    The proposed framework requires no control of the Lipschitz constant during the learning process of the denoiser. The implicit regularizer induced by the denoiser is not restricted to smooth ones (Remark 5).
  5. 7)
    Simulation results support the theoretical findings as well as showing potential advantages of the proposed approach (Section VI).
A part of Sections III, IV-A, and V-A has been presented in [44] without proofs. The present work contains many new results/discussions (such as Theorem 3 and Propositions 1 and 3) from a wider scope as well as proofs and simulations.
II Preliminaries
[…]
II-A Properness, Subdifferentiability, and Lower Semicontinuity
We consider a function $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]:={\mathbb{R}}\cup\{+\infty\}$. A function $f$ is proper if the domain is nonempty; i.e., ${\rm dom}~{}f:=\{x\in{\mathcal{H}}\mid f(x)<+\infty\}\neq\varnothing$. Given a proper function $f$, the set
${}\partial f(x):=\{z\in{\mathcal{H}}\mid\left\langle{y\!-\!x},{z}\right\rangle+f(x)\leq f(y),~{}\forall y\in{\mathcal{H}}\}$ (2)
is the subdifferential of $f$ at $x\in{\mathcal{H}}$ [39]; each element $f^{\prime}(x)\in\partial f(x)$ is a subgradient of $f$ at $x$. A function $f$ is lower semicontinuous (or closed) on ${\mathcal{H}}$ if the level set ${\rm lev}_{\leq a}f:=\left\{x\in{\mathcal{H}}:f(x)\leq a\right\}$ is closed for every $a\in{\mathbb{R}}$. Every continuous function is lower semicontinuous.
II-B Convexity and Conjugation
A function $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ is convex on ${\mathcal{H}}$ if $f(\alpha x+(1-\alpha)y)\leq\alpha f(x)+(1-\alpha)f(y)$ for every $(x,y,\alpha)\in{\rm dom}~{}f\times{\rm dom}~{}f\times[0,1]$. If the inequality of convex function holds with strict inequality whenever $x\neq y$, $f$ is strictly convex. For a positive constant $\rho\in{\mathbb{R}}_{++}$, $f$ is $\rho$-strongly convex if $f-(\rho/2)\left\|\cdot\right\|^{2}$ is convex, while $f$ is $\rho$-weakly convex if $f+(\rho/2)\left\|\cdot\right\|^{2}$ is convex.
Let $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper convex function. Then, we have (i) $\partial f(x)\neq\varnothing$ if $f$ is continuous at $x\in{\mathcal{H}}$ [39, Proposition 16.17], and (ii) $\partial f(x)=\{\nabla f(x)\}$ if $f$ is Gâteaux differentiable, where $\nabla f(x)$ is the Gâteaux gradient of $f$ at $x$555 (Gâteaux and Fréchet gradients of function) Let $U$ be an open subset of ${\mathcal{H}}$. Then, a function $f:U\rightarrow{\mathbb{R}}$ is Gâteaux differentiable at a point $x\in{\mathcal{H}}$ if there exists a vector $\nabla f(x)\in{\mathcal{H}}$ such that $\lim_{\delta\rightarrow 0}\frac{f(x+\delta h)-f(x)}{\delta}=\left\langle{\nabla f(x)},{h}\right\rangle$ for every $h\in{\mathcal{H}}$. On the other hand, a function $f:U\rightarrow{\mathbb{R}}$ is Fréchet differentiable at a point $x\in{\mathcal{H}}$ if there exists a vector $\nabla f(x)\in{\mathcal{H}}$ such that $\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)-\left\langle{\nabla f(x)},{h}\right\rangle}{\left\|h\right\|}=0$. In each case, $\nabla f(x)$ is the Gâteaux (or Fréchet) gradient of $f$ at $x$. If $f$ is Fréchet differentiable over $U$, it is also Gâteaux differentiable over $U$, and the Gâteaux and Fréchet gradients coincide. Those readers who are unfamiliar with convex analysis may consider $\nabla f$ to be the standard gradient without introducing themselves to the general notion of Gâteaux differentiability.
The set of all proper lower-semicontinuous convex functions $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ is denoted by $\Gamma_{0}({\mathcal{H}})$. Given a function $f\in{\Gamma_{0}(\mathcal{H})}$, the Fenchel conjugate of $f$ is $f^{*}:{\mathcal{H}}\rightarrow(-\infty,\infty]:x\mapsto\sup_{y\in{\mathcal{H}}}\left\langle{x},{y}\right\rangle-f(y)$, satisfying (i) $f^{*}\in{\Gamma_{0}(\mathcal{H})}$ and (ii) $u\in\partial f(x)\Leftrightarrow x\in\partial f^{*}(u)$. Given a continuous convex function $f:{\mathcal{H}}\rightarrow{\mathbb{R}}$ and a positive constant $\rho\in{\mathbb{R}}_{++}$, the interplay between smoothness and strong convexity is known [39, Theorem 18.15]: $f$ is $\rho^{-1}$-smooth if and only if the conjugate $f^{*}$ is $\rho$-strongly convex.
II-C Nonexpansiveness, Cocoercivity, and Fixed Point
In the present work, the following notions of convergence and continuity will be used.
Definition 2 (Strong/weak convergence)
Let $(x_{k})_{k\in{\mathbb{N}}}\subset{\mathcal{H}}$ be a sequence of vectors in a real Hilbert space ${\mathcal{H}}$.
  1. 1.
    The sequence $(x_{k})$ converges strongly to a point $\hat{x}\in{\mathcal{H}}$ if $\lim_{k\rightarrow\infty}\left\|x_{k}-x\right\|=0$; in symbols, $x_{k}\rightarrow\hat{x}$ (as $k\rightarrow\infty$).
  2. 2.
    The sequence $(x_{k})$ converges weakly to a point $\hat{x}\in{\mathcal{H}}$ if $\lim_{k\rightarrow\infty}\left\langle{x_{k}-\hat{x}},{y}\right\rangle=0$ for every $y\in{\mathcal{H}}$; in symbols, $x_{k}\rightharpoonup\hat{x}$ (as $k\rightarrow\infty$).
If ${\mathcal{H}}$ is finite dimensional, strong convergence coincides with weak convergence, i.e., weak convergence implies strong convergence, and vice versa [39, Fact 2.33].
Definition 3 (Lipschitz continuity)
  1. 1.
    An operator $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ is Lipschitz continuous with constant $\kappa>0$ (or $\kappa$-Lipschitz continuous for short) if for every $x,y\in{\mathcal{H}}$
    $\left\|T(x)-T(y)\right\|\leq\kappa\left\|x-y\right\|.$ (3)
  2. 2.
    An operator $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ is continuous (as a mapping from the normed space $({\mathcal{H}},\left\|\cdot\right\|)$ to itself $({\mathcal{H}},\left\|\cdot\right\|)$) if $T$ is continuous at every point $\hat{x}\in{\mathcal{H}}$, i.e., for every sequence $(x_{k})_{k\in{\mathbb{N}}}$, $x_{k}\rightarrow\hat{x}$ implies $T(x_{k})\rightarrow T(\hat{x})$.
Lipschitz continuity of $T$ implies continuity of $T$, as can readily be verified by (3).
A $1$-Lipschitz continuous operator is called nonexpansive. Given a nonexpansive operator $N:{\mathcal{H}}\rightarrow{\mathcal{H}}$, the operator $T:=(1-\alpha){\rm Id}+\alpha N$ for $\alpha\in(0,1)$ is $\alpha$-averaged nonexpansive. In particular, (1/2)-averaged operator is called firmly nonexpansive. For $\beta\in{\mathbb{R}}_{++}$, an operator $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ is $\beta$-cocoercive if $\beta T$ is firmly nonexpansive. A point $x\in{\mathcal{H}}$ is a fixed point of $T$ if $T(x)=x$. The set ${\rm Fix}(T):=\{x\in{\mathcal{H}}\mid T(x)=x\}$ of all fixed points is called the fixed point set of $T$.
We define the identity operator ${\rm Id}:{\mathcal{H}}\rightarrow{\mathcal{H}}:x\mapsto x$ on ${\mathcal{H}}$. The following fact is a part of [39, Theorem 5.15] with simplification regarding the domain of the operator.
Fact 1
Let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ be a nonexpansive operator such that ${\rm Fix}\left({T}\right)\neq\varnothing$. Define $T_{\alpha_{k}}:=(1-\alpha_{k}){\rm Id}+\alpha_{k}T$ for a real-number sequence $(\alpha_{k})_{k\in{\mathbb{N}}}\subset[0,1]$ such that $\sum_{k\in{\mathbb{N}}}\alpha_{k}(1-\alpha_{k})=+\infty$.666 It suffices that $\alpha_{k}\in[\epsilon,1-\epsilon]$ ($\forall k\in{\mathbb{N}}$) for some small $\epsilon\in(0,1/2)$. Then, given an initial point $x_{0}\in{\mathcal{H}}$, the sequence $(x_{k})_{k\in{\mathbb{N}}}$ generated by
$x_{k+1}:=T_{\alpha_{k}}(x_{k}),~{}~{}~{}k\in{\mathbb{N}},$ (4)
converges weakly to a point $\hat{x}\in{\rm Fix}\left({T}\right)$; i.e., $x_{k}\rightharpoonup\hat{x}\in{\rm Fix}\left({T}\right)$.
The update rule in (4) is called the Krasnosel’skĭ-Mann iterate. In particular, letting $\alpha_{k}:=\alpha$ $(\forall k\in{\mathbb{N}})$ for some $\alpha\in(0,1)$ in Fact 1 implies that, given any averaged nonexpansive mapping $T_{\alpha}$ with ${\rm Fix}\left({T_{\alpha}}\right)\neq\varnothing$, it holds that $T_{\alpha}^{k}(x_{0})\rightharpoonup\hat{x}\in{\rm Fix}\left({T_{\alpha}}\right)$.
II-D Monotonicity
The subdifferential operator $\partial f:{\mathcal{H}}\rightarrow 2^{\mathcal{H}}$ defined in (2) maps a vector $x$ to the set (of subgradients). In general, $T:{\mathcal{H}}\rightarrow 2^{{\mathcal{H}}}$ is called a set-valued operator, where $2^{{\mathcal{H}}}$ is the power set of ${\mathcal{H}}$; i.e., the family of all subsets of ${\mathcal{H}}$. An operator $T:{\mathcal{H}}\rightarrow 2^{{\mathcal{H}}}$ is monotone if
$\left\langle{x\!-\!y},{u\!-\!v}\right\rangle\geq 0,~{}\forall(x,u)\in{\rm gra~{}}T,~{}\forall(y,v)\in{\rm gra~{}}T,$ (5)
where gra $T:=\{(x,u)\in{\mathcal{H}}^{2}\mid u\in T(x)\}$ is the graph of $T$. A monotone operator $T:{\mathcal{H}}\rightarrow 2^{{\mathcal{H}}}$ is maximally monotone if no other monotone operator has its graph containing gra $T$ properly. For instance, the subdifferential operator $\partial f$ of a proper function $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ is monotone, and it is maximally monotone if $f\in\Gamma_{0}({\mathcal{H}})$. Suppose that a proper function $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ is Gâteaux differentiable on ${\rm dom}~{}f$ which is open and convex. Then, $f$ is convex if and only if $\nabla f$ is monotone [39, Proposition 17.7].
II-E Proximity Operator and Moreau Envelope
Motivated by Proposition 1, we define the single-valued proximity operator (the s-prox operator for short) of possibly nonconvex functions as below.
Definition 4 (Single-valued proximity (s-prox) operator)
Let $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper function. Given a positive constant $\gamma>0$, assume that $f+(1/2\gamma)\left\|x-\cdot\right\|^{2}$ has a unique minimizer for every fixed $x\in{\mathcal{H}}$. Then, the s-prox operator of $f$ of index $\gamma$ is defined by777 For proper (not necessarily convex) functions, the proximity operator is often defined as a set-valued operator because the $f+(1/2\gamma)\left\|x-\cdot\right\|^{2}$ may have multiple minimizers [45, 46] (see also https://proximity-operator.net/proximityoperator.html). In the present work, we focus on the case in which a unique minimizer exists over the entire space. Although the multi-valued proximity operator coincides with our s-prox operator when a unique minimizer exists, we give the specific name to our operator and we use the different notation ${{\rm s\mbox{-}Prox}}_{\gamma f}$ to convey our message that there are remarkable advantages from the viewpoints of explainability and optimization if we restrict ourselves to the unique-minimizer case.
${{\rm s\mbox{-}Prox}}_{\gamma f}:{\mathcal{H}}\!\rightarrow\!{\mathcal{H}}:x\mapsto\operatornamewithlimits{argmin}_{y\in\mathcal{H}}\!\Big{(}f(y)+\frac{1}{2\gamma}\left\|x\!-\!y\right\|^{2}\Big{)}.$ (6)
In the particular case of $f\in{\Gamma_{0}(\mathcal{H})}$, existence and uniqueness of minimizer is automatically ensured owing to the strict convexity of $\left\|\cdot\right\|^{2}$ and the coercivity of $\left\|\cdot\right\|^{2}$ (i.e., $\left\|y\right\|^{2}\rightarrow+\infty$ as $\left\|y\right\|\rightarrow+\infty$). In this convex case, ${{\rm s\mbox{-}Prox}}_{\gamma f}$ reduces to the classical Moreau’s proximity operator [47, 40, 48] which is firmly nonexpansive, and the following identity (known as Moreau’s decomposition) holds:
${\rm Id}={\rm s\mbox{-}Prox}_{\gamma f}+\gamma{\rm s\mbox{-}Prox}_{\gamma^{-1}f^{*}}\circ(\gamma^{-1}{\rm Id}).$ (7)
Given a function $f\in\Gamma_{0}({\mathcal{H}})$, its Moreau envelope [47, 40, 48] of index $\gamma\in(0,\infty)$ is defined as follows:
${}^{\gamma}f:{\mathcal{H}}\rightarrow{\mathbb{R}}:x\mapsto\min_{y\in{\mathcal{H}}}\Big{(}f(y)+\frac{1}{2\gamma}\left\|x-y\right\|^{2}\Big{)}.$ (8)
The Moreau envelope ${}^{\gamma}f$ is convex and Fréchet differentiable with gradient given by
$\nabla~{}^{\gamma}f(x)=\gamma^{-1}\left(x-{\rm s\mbox{-}Prox}_{\gamma f}(x)\right),$ (9)
which is $\gamma^{-1}$-Lipschitz continuous [47, 40, 48, 49]. This means that the Moreau envelope is a smooth approximation of a potentially nondifferentiable convex function $f\in{\Gamma_{0}(\mathcal{H})}$. See [49] for more details.
III Fundamental Results
The operator ${\rm s\mbox{-}Prox}_{\varphi}$ in (6) is well-defined if $\varphi:=\gamma f$ is $\rho$-weakly convex for $\rho\in(0,1)$ so that $\varphi+(1/2)\left\|x-\cdot\right\|^{2}$ is strongly (and thus strictly) convex and coercive. This simple observation is thoroughly investigated, followed by an extension of Moreau’s decomposition.
III-A Weakly Convex Function and Gradient Operator of Smooth Convex Function
We show that a given nonlinear operator $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ is the proximity operator of some weakly convex function if and only if it can be expressed as the gradient operator of a smooth convex function, i.e., a MoL-Grad denoiser (see Definition 1). The proof relies on the following lemma.
Lemma 1
For a proper function $h:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ and a vector $x\in{\mathcal{H}}$, let $J_{x}(y):=h(y)+(1/2)\left\|y-x\right\|^{2}$, $y\in{\mathcal{H}}$. Then, the following equalities hold:
$\partial J_{x}(y)=$ $~{}\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}(y)-x,$ (10)
$\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}J_{x}(y)=$ $~{}\Big{[}\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}\Big{]}^{-1}(x)$ (11)
::=:= $~{}\Big{\{}p\in{\mathcal{H}}\mid x\in\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}(p)\Big{\}}.$
In particular, if $h+(1/2)\left\|\cdot\right\|^{2}$ is coercive and strictly convex, $J_{x}$ has a unique minimizer, which is
${\rm s\mbox{-}Prox}_{h}(x)=\Big{[}\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}\Big{]}^{-1}(x)$ (12)
(in the sense of (6)).
Proof.
Fix $y\in{\mathcal{H}}$ arbitrarily. For $s\in{\mathcal{H}}$, the following equivalence holds:
$~{}s\in\partial J_{x}(y)$
$\Leftrightarrow$ $~{}\left\langle{z-y},{s}\right\rangle+J_{x}(y)\leq J_{x}(z),~{}\forall z\in{\mathcal{H}}$
$\Leftrightarrow$ $~{}\left\langle{z-y},{s}\right\rangle+h(y)+\frac{1}{2}\left\|y\right\|^{2}-\left\langle{y},{x}\right\rangle$
$\hskip 75.0001pt\leq h(z)+\frac{1}{2}\left\|z\right\|^{2}-\left\langle{z},{x}\right\rangle,~{}\forall z\in{\mathcal{H}}$
$\Leftrightarrow$ $\left\langle{z-y},{s+x}\right\rangle+h(y)+\frac{1}{2}\left\|y\right\|^{2}\leq h(z)+\frac{1}{2}\left\|z\right\|^{2}\!\!\!,~{}\forall z\in{\mathcal{H}}$
$\Leftrightarrow$ $~{}s\in\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}(y)-x,$
which verifies (10). Now, let $p\in{\mathcal{H}}$. Then, from (10), the following equivalence holds:
$p\in\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}J_{x}(y)\Leftrightarrow 0\in\partial J_{x}(p)=\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}(p)-x$
$\Leftrightarrow x\in\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}(p)\Leftrightarrow p\in\Big{[}\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}\Big{]}^{-1}(x)$ ,
which verifies (11).
When $h+(1/2)\left\|\cdot\right\|^{2}$ is coercive and strictly convex, the coercivity and strict convexity of $J_{x}$ ensures the existence and uniqueness of its minimizer, respectively, and thus (11), together with the definition of the proximity operator in (6), implies (12). ∎
Theorem 1
Let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$. Then, for every $\beta\in(0,1)$, the following two conditions are equivalent.
  1. (C1)
    $T={\rm s\mbox{-}Prox}_{\varphi}$ for some $\varphi:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ such that $\varphi+\left((1-\beta)/2\right)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$.
  2. (C2)
    $T$ is a MoL-Grad denoiser (see Definition 1), i.e., the following hold jointly.
    1. 1)
      $T=\nabla\psi$ for some Fréchet differentiable convex function $\psi\in{\Gamma_{0}(\mathcal{H})}$.
    2. 2)
      $T$ is $\beta$-cocoercive, or equivalently $\beta^{-1}$-Lipschitz continuous (by the Baillon Haddad theorem [50]).
In particular, the following statements hold.
  1. (a)
    Assume that (C1) is satisfied. Define $\check{\varphi}:=\varphi+(\left(1-\beta\right)/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$. Then, it holds that
    $T={\rm s\mbox{-}Prox}_{\varphi}=\nabla\underbrace{\Big{(}\varphi+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}^{*}}_{=\psi}=\nabla~{}^{\beta}(\check{\varphi}^{*}),$ (13)
    which is $\beta$-cocoercive (thus $\beta^{-1}$-Lipschitz continuous and maximally monotone).
  2. (b)
    Assume that (C2) is satisfied. Then, it holds that
    $T=\nabla\psi={\rm s\mbox{-}Prox}_{\psi^{*}-(1/2)\left\|\cdot\right\|^{2}},$ (14)
    where
    $\varphi=\psi^{*}-\frac{1}{2}\left\|\cdot\right\|^{2}$ (15)
    is $(1-\beta)$-weakly convex.
Proof.
C1 $\Rightarrow$ C2: For the sake of readability, we present a proof which the reader may follow directly with a minimal use of the known results in convex analysis.888 The proof can be shortened by using [39, Example 22.7] and [39, Example 14.1] to verify (16) and (18). Owing to the convexity of $\check{\varphi}:=\varphi+((1-\beta)/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$, we have $T^{-1}={\rm s\mbox{-}Prox}_{\varphi}^{-1}=\partial(\varphi+(1/2)\left\|\cdot\right\|^{2})=\partial(\check{\varphi}+(\beta/2)\left\|\cdot\right\|^{2})=\partial\check{\varphi}+\beta{\rm Id}$ by Lemma 1. Given any fixed $x,y\in{\mathcal{H}}$, let $u\in T^{-1}(x)$ and $v\in T^{-1}(y)$ ($\Leftrightarrow x=T(u),~{}y=T(v)$). Then, since $T^{-1}-\beta{\rm Id}=\partial\check{\varphi}$ is a monotone operator, it follows that
$\left\langle{x-y},{(u-\beta x)-(v-\beta y)}\right\rangle\geq 0$
$\Leftrightarrow$ $\left\langle{x-y},{u-v}\right\rangle\geq\beta\left\|x-y\right\|^{2}$
$\Leftrightarrow$ $\left\langle{T(u)-T(v)},{u-v}\right\rangle\geq\beta\left\|T(u)-T(v)\right\|^{2}$
$\Leftrightarrow$ $\left\langle{\beta T(u)-\beta T(v)},{u-v}\right\rangle\geq\left\|\beta T(u)-\beta T(v)\right\|^{2},$ (16)
where the last inequality implies that $T$ is $\beta$-cocoercive. On the other hand, for every $y\in{\mathcal{H}}$, it holds that
$[^{\beta}(\check{\varphi}^{*})]^{*}(y)=$ $\sup_{x\in{\mathcal{H}}}\Big{[}\left\langle{x},{y}\right\rangle-\inf_{z\in{\mathcal{H}}}\Big{(}\check{\varphi}^{*}(z)+\frac{1}{2\beta}\left\|z-x\right\|^{2}\Big{)}\Big{]}$
$=$ $\sup_{x\in{\mathcal{H}}}\Big{[}\!\left\langle{x},{y}\right\rangle+\sup_{z\in{\mathcal{H}}}\Big{(}\!\!-\check{\varphi}^{*}(z)-\frac{1}{2\beta}\left\|z-x\right\|^{2}\Big{)}\Big{]}$
$=$ $\sup_{z\in{\mathcal{H}}}\Big{[}\!-\check{\varphi}^{*}(z)+\sup_{x\in{\mathcal{H}}}\Big{(}\!\left\langle{x},{y}\right\rangle-\frac{1}{2\beta}\left\|z-x\right\|^{2}\!\Big{)}\Big{]}$
$=$ $\sup_{z\in{\mathcal{H}}}\Big{[}-\check{\varphi}^{*}(z)+\left\langle{z},{y}\right\rangle+\frac{\beta}{2}\left\|y\right\|^{2}\Big{]}$
$=$ $~{}\check{\varphi}^{**}(y)+\frac{\beta}{2}\left\|y\right\|^{2}=\check{\varphi}(y)+\frac{\beta}{2}\left\|y\right\|^{2},$ (17)
where $\check{\varphi}^{**}=\check{\varphi}$ because $\check{\varphi}\in\Gamma_{0}({\mathcal{H}})$ by virtue of the Fenchel-Moreau theorem [39, Theorem 13.37]. From (17), it follows that
$\Big{(}\check{\varphi}+\frac{\beta}{2}\left\|\cdot\right\|^{2}\Big{)}^{*}=[^{\beta}(\check{\varphi}^{*})]^{**}=~{}^{\beta}(\check{\varphi}^{*})=:\psi,$ (18)
where (${}^{\beta}(\check{\varphi}^{*})$, and thus $\psi$, is a Fréchet differentiable function with $\beta^{-1}$-Lipschitz continuous gradient (see Section II-E):
$\nabla\psi=\nabla\Big{(}\check{\varphi}+\frac{\beta}{2}\left\|\cdot\right\|^{2}\Big{)}^{*}=\Big{[}\partial\Big{(}\check{\varphi}+\frac{\beta}{2}\left\|\cdot\right\|^{2}\Big{)}\Big{]}^{-1}=T.$ (19)
The proof in this part also verifies Theorem 1(a).
C2 $\Rightarrow$ C1: We observe that $T^{-1}=(\nabla\psi)^{-1}=\partial\psi^{*}=\partial(\varphi+(1/2)\left\|\cdot\right\|^{2})$, where $\varphi:=\psi^{*}-(1/2)\left\|\cdot\right\|^{2}$. This leads to $T=[\partial(\varphi+(1/2)\left\|\cdot\right\|^{2})]^{-1}={\rm s\mbox{-}Prox}_{\varphi}$. Here, the Fréchet differentiability of $\psi$ and the $\beta^{-1}$-Lipschitz continuity of $\nabla\psi$ imply $\beta$-strong convexity of the conjugate $\psi^{*}$ (see Section II-B), and we thus obtain $\varphi+\left((1-\beta)/2\right)\left\|\cdot\right\|^{2}=\psi^{*}-(\beta/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$. The proof in this part also verifies Theorem 1(b). ∎
The relations among $\varphi$, $\psi$, and $\check{\varphi}$ are depicted in Fig. 1. An application of Theorem 1(b) has been shown in [51], where an external division of two Moreau’s proximity operators is shown to be the proximity operator of a weakly convex function under a certain condition (see Example 4 in Section V-B).
[…]
Figure 1: A weakly convex function $\varphi$, a smooth function $\psi$, and the intermediate “neutrally” convex function $\check{\varphi}$ between $\varphi$ and the strongly convex function $\psi^{*}$.
Remark 1 (Alternative proof of Theorem 1(b))
Theorem 1(b) could be proved using known results as follows. Since $\nabla\psi$ is $\beta^{-1}$-Lipschitz continuous under condition (C2), it follows by [39, Theorem 18.15] that $\psi={}^{\beta}(\check{\varphi}^{*})$, where $\check{\varphi}:=\psi^{*}-(\beta/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$. Hence, by Moreau’s decomposition, it follows that $\nabla\psi=\beta^{-1}({\rm Id}-{\rm s\mbox{-}Prox}_{\beta\check{\varphi}^{*}})={\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ(\beta^{-1}{\rm Id})={\rm s\mbox{-}Prox}_{\beta^{-1}\psi^{*}-(1/2)\left\|\cdot\right\|^{2}}\circ(\beta^{-1}{\rm Id})={\rm s\mbox{-}Prox}_{\psi^{*}-(1/2)\left\|\cdot\right\|^{2}},$ where the last inequality can be verified by the definition of proximity operator with some manipulations. Here, it is clear that $\varphi:=\psi^{*}-(1/2)\left\|\cdot\right\|^{2}=\check{\varphi}-((1-\beta)/2)\left\|\cdot\right\|^{2}$ is $(1-\beta)$-weakly convex.
III-B Extension of Moreau’s Decomposition
We show that Moreau’s decomposition (see Section II-E) can be extended to weakly convex functions.
Proposition 2
Let $\varphi:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper function such that $\check{\varphi}:=\varphi+((1-\beta)/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ for $\beta\in(0,1)$. Then, the following equalities hold:
${\rm s\mbox{-}Prox}_{\varphi}$ $={\rm s\mbox{-}Prox}_{\check{\varphi}/\beta}\circ\beta^{-1}{\rm Id},$ (20)
${\rm Id}$ $=\beta{\rm s\mbox{-}Prox}_{\varphi}+{\rm s\mbox{-}Prox}_{\beta\check{\varphi}^{*}}.$ (21)
Proof.
By using Moreau’s decomposition, it can be verified that
$\beta^{-1}\left({\rm Id}-{\rm s\mbox{-}Prox}_{\beta\check{\varphi}^{*}}\right)(x)={\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}(\beta^{-1}x)$
$=\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}\left(\beta^{-1}\check{\varphi}(y)+\frac{1}{2}\left\|y-\beta^{-1}x\right\|^{2}\right)$
$=\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}\left(\varphi(y)+\frac{1-\beta}{2}\left\|y\right\|^{2}+\frac{\beta}{2}\left\|y-\beta^{-1}x\right\|^{2}\right)$
$=\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}\left(\varphi(y)+\frac{1}{2}\left\|y-x\right\|^{2}+K_{x}\right)={\rm s\mbox{-}Prox}_{\varphi}(x),$
where $K_{x}:=((\beta^{-1}-1)/2)\left\|x\right\|^{2}$ is constant in $y$. ∎
The identity in (20) can be written equivalently as follows:
$\underbrace{{\rm s\mbox{-}Prox}_{\check{\varphi}/\beta}}_{\mbox{firmly nonexpansive}}=\underbrace{{\rm s\mbox{-}Prox}_{\varphi}}_{\beta\mbox{-cocoercive}}\circ\beta{\rm Id}.$ (22)
The relation in (22) is consistent with Theorem 1 which shows $\beta$-cocoercivity of ${\rm s\mbox{-}Prox}_{\varphi}$ (see Lemma B.1 in Appendix B).
Remark 2 (On Proposition 2)
As opposed to the case of ${\rm s\mbox{-}Prox}_{\varphi}$ for the weakly convex function $\varphi$, ${\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}$ and ${\rm s\mbox{-}Prox}_{\beta\check{\varphi}^{*}}$ are the proximity operators in the sense of Moreau, because $\beta^{-1}\check{\varphi}$ and $\beta\check{\varphi}^{*}$ are convex. The identity in (20) gives an expression of the s-prox operator of the weakly convex function in terms of the “classical” proximity operator of the convex function $\beta^{-1}\check{\varphi}$. While the identity in (21) resembles Moreau’s decomposition, the essential difference is that $\varphi$ is weakly convex. More specifically, (21) gives a way of decomposing a vector using the proximity operators of the weakly convex function $\varphi$ and the convex function $\check{\varphi}^{*}$.
Proposition 2 can be given in a general form as below.
Proposition 3
Let $\phi:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ such that $\check{\phi}:=\phi+(\rho/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ for a positive constant $\rho\in{\mathbb{R}}_{++}$. Then, for every positive constant $\sigma\in{\mathbb{R}}_{++}$, it holds that
${\rm s\mbox{-}Prox}_{\check{\phi}/\sigma}={\rm s\mbox{-}Prox}_{\phi/(\sigma+\rho)}\circ\left(\frac{\sigma}{\sigma+\rho}{\rm Id}\right),$ (23)
${\rm Id}={\rm s\mbox{-}Prox}_{\sigma\check{\phi}^{*}}+\sigma{\rm s\mbox{-}Prox}_{\phi/(\sigma+\rho)}\circ(\sigma+\rho)^{-1}{\rm Id}.$ (24)
Proof.
Let $\varphi:=\phi/(\sigma+\rho)$ in (22). Then, $\varphi$ is $(1-\beta)$-weakly convex for $\beta:=\sigma/(\sigma+\rho)\in(0,1)$, and we thus obtain (23). Using Moreau’s decomposition (7) and then applying (23) yield
${\rm Id}=$ $~{}{\rm s\mbox{-}Prox}_{\sigma\check{\phi}^{*}}+\sigma{\rm s\mbox{-}Prox}_{\check{\phi}/\sigma}\circ(\sigma^{-1}{\rm Id})$
$=$ $~{}{\rm s\mbox{-}Prox}_{\sigma\check{\phi}^{*}}+\sigma{\rm s\mbox{-}Prox}_{\phi/(\sigma+\rho)}\circ(\sigma+\rho)^{-1}{\rm Id}.$ (25)
IV Operator-Regularization Approaches — Characterization of Implicit Regularizer and Convergence to Optimal Point
We have seen that a MoL-Grad denoiser can always be expressed as $T={\rm s\mbox{-}Prox}_{\varphi}$ for the $(1-\beta)$-weakly convex function $\varphi=\psi^{*}-(1/2)\left\|\cdot\right\|^{2}$ (Theorem 1). We use this fundamental result to make the plug-and-play methods “transparent (a while box)”. More specifically, we consider operator splitting algorithms with the proximity operators replaced by our MoL-Grad denoiser $T$, which accommodates prior information (often implicitly).
We shall show that such plug-and-play methods (employing our denoiser $T=\nabla\psi$) asymptotically optimize a certain cost function in the sense of generating a vector sequence convergent to its minimizer. Here, the cost function involves an implicit “operator-inducing” regularizer which depends on $\varphi=\psi^{*}-(1/2)\left\|\cdot\right\|^{2}$. In this way, the loss function is implicitly regularized by the denoiser $T$, as opposed to the case of the traditional functional-regularization approaches which explicitly regularize the loss by some penalty functions.
In the following, two specific operator-regularization algorithms are studied. A generic way of applying the same idea systematically to different algorithms is then presented.
IV-A Forward-Backward Splitting Type Algorithm
Based on the results given in Section III, we show that the algorithm in (1) with a MoL-Grad denoiser $T$ converges to a solution of an optimization problem.
Theorem 2
Let $f\in{\Gamma_{0}(\mathcal{H})}$ be a $\kappa$-smooth $\rho$-strongly-convex function with $\kappa>\rho>0$. Assume that $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ is a MoL-Grad denoiser (see Definition 1) for $\beta\in((\kappa-\rho)/(\kappa+\rho),1)\subsetneq(0,1)$ so that the Lipschitz constant is bounded by $\beta^{-1}<(\kappa+\rho)/(\kappa-\rho)$. Set $\mu\in[(1-\beta)/\rho,(1+\beta)/\kappa)$. Then, the following hold.
  1. 1.
    Let $\hat{f}:=f-[(1-\beta)/(2\mu)]\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ and $\check{\varphi}:=\varphi+[(1-\beta)/2]\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ so that $\mu\hat{f}+\check{\varphi}=\mu f+\varphi$. Then, it holds that
    $T\circ({\rm Id}-\mu\nabla f)={\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ({\rm Id}-\beta^{-1}\mu\nabla\hat{f})$ (26)
    with $\beta^{-1}\in(1,2/L_{\mu\nabla\hat{f}})$, where $L_{\mu\nabla\hat{f}}:=\mu\kappa-(1-\beta)>0$ is the Lipschitz constant of $\mu\nabla\hat{f}$.
  2. 2.
    Suppose that $\mu f+\varphi$ has a minimizer in ${\mathcal{H}}$. Then, for an arbitrary $x_{0}\in{\mathcal{H}}$, the sequence $(x_{k})_{k\in{\mathbb{N}}}\subset{\mathcal{H}}$ generated by (1) converges weakly to a minimizer $\hat{x}$ of $\mu f+\varphi$. (In this case, $\varphi$ is the implicit regularizer.)
Proof.
1) Since $\mu\geq(1-\beta)/\rho\Leftrightarrow\rho\geq(1-\beta)/\mu$, the $\rho$-strong convexity of $f$ immediately implies $\hat{f}\in\Gamma_{0}({\mathcal{H}})$. On the other hand, since $\varphi$ is $(1-\beta)$-weakly convex by Theorem 1, $\hat{\varphi}\in\Gamma_{0}({\mathcal{H}})$ follows. Now, using Theorem 1(b) and Proposition 2, we can verify that
$T\circ({\rm Id}-\mu\nabla f)$ $={\rm s\mbox{-}Prox}_{\varphi}\circ({\rm Id}-\mu\nabla f)$
$={\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ(\beta^{-1}{\rm Id}-\beta^{-1}\mu\nabla f)$
$={\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ({\rm Id}-\beta^{-1}\mu\nabla\hat{f}).$
It can be verified that $L_{\mu\nabla\hat{f}}=\mu\kappa-(1-\beta)>\mu\rho-(1-\beta)\geq 0$, where the assumption $\mu\geq(1-\beta)/\rho$ is used again to verify the last inequality. On the other hand, we observe that $\mu<(1+\beta)/\kappa\Leftrightarrow\mu\kappa-(1-\beta)<2\beta\Leftrightarrow\beta^{-1}<2/(\mu\kappa-(1-\beta))=2/L_{\mu\nabla\hat{f}}$, and thus $\beta^{-1}<2/L_{\mu\nabla\hat{f}}$ is verified. Finally, since $\nabla f$ is $\kappa$-Lipschitz continuous by the assumption, it is not difficult to see that $\mu\nabla\hat{f}=\mu\nabla f-(1-\beta){\rm Id}$ is $(\mu\kappa-(1-\beta))$-Lipschitz continuous (see, e.g., [52, Lemma 8]).
2) By Theorem 2.1, (1) can be regarded as the forward-backward splitting iterate for the sum of the convex functions $\mu\hat{f}$ and $\check{\varphi}$ with the step size $\beta^{-1}\in(1,2/L_{\mu\nabla\hat{f}})$. Hence, from the standard argument of the forward-backward splitting method (see, e.g., [49, Example 17.6]), the composition ${\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ({\rm Id}-\beta^{-1}\mu\nabla\hat{f})$ is an averaged nonexpansive operator with the fixed-point set ${\rm Fix}\left({{\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ({\rm Id}-\beta^{-1}\mu\nabla\hat{f})}\right)=\operatornamewithlimits{argmin}_{x\in{\mathcal{H}}}[\mu\hat{f}(x)+\check{\varphi}(x)]=\operatornamewithlimits{argmin}_{x\in{\mathcal{H}}}[\mu f(x)+\varphi(x)]\neq\varnothing$, where the nonemptiness is due to the assumption. The classical Krasnosel’skĭ-Mann iterate (see Fact 1) thus ensures weak convergence of $(x_{k})_{k\in{\mathbb{N}}}$ to a fixed point which is a minimizer of $\mu f(x)+\varphi(x)$. ∎
The existence (and uniqueness) of minimizer of $\mu f+\varphi$ is ensured if $\mu\rho>1-\beta$. The case in which $f$ is not strongly convex has been addressed in [51]. Note in Theorem 2.2 that $(x_{k})_{k\in{\mathbb{N}}}$ converges “strongly (in the norm sense)” to $\hat{x}$ if ${\mathcal{H}}$ is finite dimensional (see Definition 2).
Remark 3 (On Theorem 2)
If $\beta\approx 1$ (i.e., $T$ is nearly nonexpansive), the range of $\mu$ is approximately identical to $(0,2/\kappa)$. Meanwhile, if $\kappa\approx\rho$, the range of $\beta$ is nearly $(0,1)$, meaning that the Lipschitz constant $\beta^{-1}$ of $T$ could be an arbitrary real number virtually. However, focusing on the gap with respect to the range of $\mu$, we observe that $(1+\beta)/\kappa-(1-\beta)/\rho=K_{\kappa,\rho}(\beta(\kappa+\rho)/(\kappa-\rho)-1)>0$ with $K_{\kappa,\rho}:=\rho^{-1}-\kappa^{-1}>0$, where the strict positivity of the gap is verified by $\beta>(\kappa-\rho)/(\kappa+\rho)$. This suggests that the gap of $\mu$ vanishes when $\beta^{-1}$ is significantly large so that $\beta\approx(\kappa-\rho)/(\kappa+\rho)$.
IV-B Primal-Dual Splitting Type Algorithm
We now build a framework which could be applied to a wider class of problems involving total variation, for instance, where a bounded linear operator $L:{\mathcal{H}}\rightarrow\mathcal{U}$ will be used in the algorithm. Here, ${\mathcal{H}}$ and $\mathcal{U}$ are real Hilbert spaces equipped with the inner products $\left\langle{\cdot},{\cdot}\right\rangle_{{\mathcal{H}}}$, $\left\langle{\cdot},{\cdot}\right\rangle_{\mathcal{U}}$, and the induced norms $\left\|\cdot\right\|_{{\mathcal{H}}}$, $\left\|\cdot\right\|_{\mathcal{U}}$, respectively. We define the operator norm $\left\|L\right\|:=\sup_{x\in{\mathcal{H}},\left\|x\right\|_{{\mathcal{H}}}\leq 1}\left\|Lx\right\|_{\mathcal{U}}$. The adjoint operator of $L$ is denoted by $L^{*}:\mathcal{U}\rightarrow{\mathcal{H}}$. With a nonlinear operator $T:\mathcal{U}\rightarrow\mathcal{U}$, we consider the following algorithm inspired by the Condat–Vũ algorithm [8, 9].999 An earlier study of an operator-regularization approach using the primal-dual splitting algorithm can be found in [53].
Algorithm 1
Initialization: $x_{0}\in{\mathcal{H}}$, $u_{0}\in\mathcal{U}$
Requirements: $\sigma>0$, $\tau>0$, $\rho>0$
$\tilde{u}_{k+1}:=u_{k}+\sigma Lx_{k}$
$u_{k+1}:=\tilde{u}_{k+1}-\sigma T\bigg{(}\bigg{(}\sigma+\dfrac{\rho}{\left\|L\right\|^{2}}\bigg{)}^{-1}\tilde{u}_{k+1}\bigg{)}$
$x_{k+1}:=\bigg{(}{\rm Id}+\dfrac{\tau\rho}{\left\|L\right\|^{2}}L^{*}L\bigg{)}x_{k}-\tau\nabla f(x_{k})-\tau L^{*}(2u_{k+1}-u_{k})$
[…]
Figure 2: A schematic diagram of the derivation of Algorithm 1.
Theorem 3
Let $f\in{\Gamma_{0}(\mathcal{H})}$ be a smooth $\rho$-strongly convex function for $\rho\in{\mathbb{R}}_{++}$ so that the function
$\Gamma_{0}({\mathcal{H}})\ni\hat{f}:x\mapsto f(x)-\frac{\rho}{2\left\|L\right\|^{2}}\left\|Lx\right\|_{\mathcal{U}}^{2}$ (27)
is $\kappa$-smooth for some $\kappa\in{\mathbb{R}}_{++}$. Assume that $T:\mathcal{U}\rightarrow\mathcal{U}$ is a MoL-Grad denoiser for some $\psi\in\Gamma_{0}({\mathcal{H}})$ and $\beta\in(0,1)$. Set the step sizes $\sigma,\tau>0$ such that
  • (i)
    $\sigma\leq\dfrac{\rho\beta}{\left\|L\right\|^{2}(1-\beta)}$,
  • (ii)
    $\tau\Big{(}\sigma\left\|L\right\|^{2}+\dfrac{\kappa}{2}\Big{)}<1$.
Then, for arbitrary $(x_{0},u_{0})\in{\mathcal{H}}\times\mathcal{U}$, the sequences $(x_{k})_{k\in{\mathbb{N}}}\subset{\mathcal{H}}$ and $(u_{k})_{k\in{\mathbb{N}}}\subset\mathcal{U}$ generated by Algorithm 1 converge weakly to solutions $\hat{x}\in{\mathcal{H}}$ and $\hat{u}\in\mathcal{U}$ of the following primal and dual problems, respectively:
$\min_{x\in{\mathcal{H}}}f(x)+g(Lx),$ (28)
$\min_{u\in\mathcal{U}}\hat{f}^{*}(-L^{*}u)+\check{g}^{*}(u),$ (29)
provided that such solutions exist. Here,
$g:=$ $~{}\Big{(}\sigma+\frac{\rho}{\left\|L\right\|^{2}}\Big{)}\varphi=\Big{(}\sigma+\frac{\rho}{\left\|L\right\|^{2}}\Big{)}\Big{(}\psi^{*}-\frac{1}{2}\left\|\cdot\right\|_{\mathcal{U}}^{2}\Big{)},$ (30)
$\check{g}:=$ $~{}g+\frac{\rho}{2\left\|L\right\|^{2}}\left\|\cdot\right\|_{\mathcal{U}}^{2}\in\Gamma_{0}(\mathcal{U})$ (31)
satisfy $f+g\circ L=\hat{f}+\check{g}\circ L\in\Gamma_{0}({\mathcal{H}})$. (In this case, $g\circ L$ is the implicit regularizer.)
Proof.
First of all, ${\mathcal{H}}\ni x\mapsto(\rho/2)\left\|x\right\|_{{\mathcal{H}}}^{2}-(\rho/(2\left\|L\right\|^{2}))\left\|Lx\right\|_{\mathcal{U}}^{2}$ is a convex function owing to the positive semidefiniteness of ${\rm Id}-L^{*}L/\left\|L\right\|^{2}$. This, together with the $\rho$-strong convexity of $f$, implies that $\hat{f}=f-(\rho/(2\left\|L\right\|^{2}))\left\|L\cdot\right\|_{\mathcal{U}}^{2}$ is convex.
Since $\varphi$ is $(1-\beta)$-weakly convex by Theorem 1, $g=(\sigma+\rho/\left\|L\right\|^{2})\varphi$ is $(\sigma+\rho/\left\|L\right\|^{2})(1-\beta)$-weakly convex. On the other hand, condition (i) of the theorem implies that $(\sigma+\rho/\left\|L\right\|^{2})(1-\beta)\leq\rho/\left\|L\right\|^{2}$, and hence $\check{g}$ is convex as well.
We show that Algorithm 1 is obtained by applying the Condat–Vũ algorithm to the primal and dual problems, (28) and (29). By $\nabla\hat{f}(x_{k})=\nabla f(x_{k})-(\rho/\left\|L\right\|^{2})L^{*}Lx_{k}$, the last recursion of Algorithm 1 can be rewritten as
$x_{k+1}=x_{k}-\tau\nabla\hat{f}(x_{k})-\tau L^{*}(2u_{k+1}-u_{k}).$ (32)
We shall now show that the first two recursions can be expressed as
$u_{k+1}={\rm s\mbox{-}Prox}_{\sigma\check{g}^{*}}(u_{k}+\sigma Lx_{k}),$ (33)
because the set of recursions (32) and (33) gives (a no-relaxation version of) the Condat–Vũ iteration (form II) [11]. Indeed, the formulation in [11] involves three functions $f$, $g$, and $h$, each of which corresponds to the zero function, $\check{g}$, and $\hat{f}$, respectively, in our case. Once (33) is proved, the convergence follows immediately by [11, Theorem 7.1] under condition (ii).
Letting $\phi:=g$ in Proposition 3 yields
${\rm s\mbox{-}Prox}_{\sigma\check{g}^{*}}={\rm Id}-\sigma{\rm s\mbox{-}Prox}_{g/(\sigma+\rho/\left\|L\right\|^{2})}\circ\bigg{(}\sigma+\frac{\rho}{\left\|L\right\|^{2}}\bigg{)}^{-1}{\rm Id},$ (34)
which, together with $T={\rm s\mbox{-}Prox}_{\varphi}={\rm s\mbox{-}Prox}_{g/(\sigma+\rho/\left\|L\right\|^{2})}$ by Theorem 1, verifies
${\rm s\mbox{-}Prox}_{\sigma\check{g}^{*}}={\rm Id}-\sigma T\circ\bigg{(}\sigma+\frac{\rho}{\left\|L\right\|^{2}}\bigg{)}^{-1}{\rm Id}.$ (35)
The identity in (35) justifies the expression in (33), and this completes the proof. ∎
Remark 4 (Parameter design)
From (30), the optimization problem (28) which Algorithm 1 actually solves depends on the dual step-size $\sigma$ as well as $\varphi$ which is associated with $T$. Note here that $\rho$ and $L$ are assumed to be given. Specifically, the parameter $\sigma$ governs the weight of the regularizer (prior) $\varphi$ relative to $f$, and a tuning parameter $\delta\in(0,1]$ can be introduced to control the weight as $\sigma:=\delta\rho\beta/[\left\|L\right\|^{2}(1-\beta)]$. It should be remarked that changing the tuning parameter(s) in $T$ (if any) may scale $\varphi$ as well as $\beta$ which determines the upper bound of $\sigma$ (see condition (i) of Theorem 3). The primal step size $\tau$ only affects the convergence speed; it can simply be set to $\tau:=\gamma(\sigma\left\|L\right\|^{2}+\kappa/2)^{-1}$, where $\gamma\in(0,1)$ (our recommended choice is $\gamma:=0.9$).
IV-C A Systematic Way of Building Operator-Regularization Algorithms using MoL-Grad with Convergence Guarantee
We have seen in Section IV-B that Algorithm 1 minimizes $f+g\circ L$ (and $\hat{f}^{*}\circ(-L^{*})+\check{g}^{*}$ at the same time) under the conditions. But, how can we derive such an algorithm? The derivation process of Algorithm 1 is illustrated in Fig. 2, elaborated below in a step-by-step manner.
Suppose that we would like to plug a (well-performing) MoL-Grad denoiser $T$ into the primal-dual splitting algorithm. Since such a $T$ is the proximity operator of a weakly convex function (Theorem 1), we start with the convex objective $f+g\circ L$ with $g$ weakly convex. First, we define $\hat{f}$ and $\check{g}$ as in (27) and (31), respectively, which satisfy $f+g\circ L=\hat{f}+\check{g}\circ L$. Second, the primal-dual algorithm is applied to the sum $\hat{f}+\check{g}\circ L$ of convex functions to obtain the set of recursions (32) and (33). Third, Proposition 3 is applied to ${\rm s\mbox{-}Prox}_{\sigma\check{g}^{*}}$ to obtain (34), and then ${\rm s\mbox{-}Prox}_{g/(\sigma+\rho/\left\|L\right\|^{2})}$ is replaced by $T$ to obtain (35). Note here that ${\rm s\mbox{-}Prox}_{g/(\sigma+\rho/\left\|L\right\|^{2})}$ is a MoL-Grad denoiser from Theorem 1 owing to the $(1-\beta)$-weak convexity of $g/(\sigma+\rho/\left\|L\right\|^{2})(=\varphi)$. Finally, plugging (35) into (33) yields Algorithm 1.
The point is that the denoiser can be expressed as $T={\rm s\mbox{-}Prox}_{g/(\sigma+\rho/\left\|L\right\|^{2})}$ based on Theorem 1, which is further linked to the (Moreau’s) proximity operator by Proposition 3. Thanks to this link, Algorithm 1 has been shown to minimize $f+g\circ L$ with the operator-inducing function $g$ given in (30). In general, given an arbitrary operator splitting algorithm, the proximity operator(s) employed can be replaced systematically by any MoL-Grad denoiser(s) by the following steps.
Step 0. Suppose that we are given an operator splitting algorithm in which the proximity operator wants to be replaced by a MoL-Grad denoiser $T$. The associated optimization problem typically involves a smooth function $f$ and a prox-friendly convex function $g$ (or possibly multiple such functions) which is possibly composed with a bounded linear operator. As opposed to this standard setting, we suppose that $g$ is weakly convex (and possibly other prox-friendly functions could also be so) while $f$ is strongly convex so that the whole objective $f+g\circ L$ is convex.
Step 1. Add a quadratic function to the weakly convex function, and subtract its corresponding quadratic function from the strongly convex function in such a way that (a) the sum remains the same, (b) the weakly convex function is convexified, and (c) the strongly convex function remains convex. By doing so, a convex optimization problem is obtained with the same objective.
Step 2. Apply the operator splitting algorithm to the convex optimization problem obtained in step 1.
Step 3. Apply Proposition 3 (or Proposition 2), and replace the proximity operator of $g$ (of some index) by the $T$ (and also replace those of the other functions by other MoL-Grad denoisers).
Through the above steps, the obtained algorithm involves no proximity operator anymore in an explicit form. By the same principle as stated above for Algorithm 1, however, it iteratively solves an optimization problem involving an implicit regularizer(s) which is induced by the denoiser $T$.
V Examples and Relation to Prior Work
We present examples of MoL-Grad denoiser, or equivalently the proximity operator of weakly convex function in the sense of Definition 4, and we then discuss the relation to prior work. In the Euclidean space setting, we use the standard notation $\left\|\cdot\right\|_{1}$ and $\left\|\cdot\right\|_{2}$ for the $\ell_{1}$ and $\ell_{2}$ norms of vector, respectively, and $(\cdot)^{\top}$ for matrix/vector transposition.
V-A Shrinkage Operators for Scalar
Below are a selection of typical shrinkage operators.
V-A1 Soft shrinkage
The soft shrinkage operator ${\rm soft}_{\lambda}:{\mathbb{R}}\rightarrow{\mathbb{R}}$ for the threshold $\lambda\in{\mathbb{R}}_{++}$ defined by
${\rm soft}(x):=\begin{cases}0&\mbox{if }\left|x\right|\leq\lambda\\ x-\lambda&\mbox{if }x>\lambda\\ x+\lambda&\mbox{if }x<-\lambda\end{cases}$ (36)
is the proximity operator of the convex function $\lambda\left|\cdot\right|$. The soft shrinkage operator is a MoL-Grad denoiser for an arbitrary $\beta\in(0,1)$ because it is (firmly) nonexpansive.
V-A2 Firm shrinkage
The firm shrinkage operator ${\rm firm}_{\lambda_{1},\lambda_{2}}:{\mathbb{R}}\rightarrow{\mathbb{R}}$ for the thresholds $\lambda_{1}\in{\mathbb{R}}_{++}$ and $\lambda_{2}\in(\lambda_{1},+\infty)$ defined by [54]
${\rm firm}_{\lambda_{1},\lambda_{2}}(x)\!:=\!\begin{cases}0&\!\!\mbox{if }\left|x\right|\leq\lambda_{1}\\ {\rm sign}(x)\frac{\lambda_{2}(\left|x\right|-\lambda_{1})}{\lambda_{2}-\lambda_{1}}&\!\!\mbox{if }\lambda_{1}<\left|x\right|\leq\lambda_{2}\\ x&\!\!\mbox{if }\left|x\right|>\lambda_{2}\end{cases}$ (37)
is the proximity operator ${\rm s\mbox{-}Prox}_{\varphi}$ of the following $\lambda_{1}/\lambda_{2}$-weakly convex function $\varphi:=\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}$, where
$\varphi_{\lambda_{2}}^{\rm MC}(x):=\begin{cases}\left|x\right|-\frac{1}{2\lambda_{2}}x^{2}&\mbox{if }\left|x\right|\leq\lambda_{2}\\ \frac{1}{2}\lambda_{2}&\mbox{if }\left|x\right|>\lambda_{2}\end{cases}$ (38)
is the minimax concave (MC) penalty [17, 52, 29].
V-A3 Garrote shrinkage
The garrote shrinkage (nonnegative garrote thresholding) operator ${\rm garrote}_{\lambda}:{\mathbb{R}}\rightarrow{\mathbb{R}}$ for the threshold $\lambda\in{\mathbb{R}}_{++}$ defined by [55]
${\rm garrote}_{\lambda}(x):=\begin{cases}0&\mbox{if }\left|x\right|\leq\lambda\\ x-\frac{\lambda^{2}}{x}&\mbox{if }\left|x\right|>\lambda\end{cases}$ (39)
is the proximity operator ${\rm s\mbox{-}Prox}_{\varphi}$ of the following 1/2-weakly convex function:
$\varphi(x):=$ $~{}\frac{1}{4}\left(\left|x\right|\sqrt{x^{2}+4\lambda^{2}}-x^{2}\right)$
$~{}+\lambda^{2}\left[\log(\left|x\right|+\sqrt{x^{2}+4\lambda^{2}})-\log 2\lambda\right].$ (40)
Note that the hard shrinkage operator is not the proximity operator of any weakly convex function (in the sense of Definition 4) because of its discontinuity.
V-B Shrinkage Operators for Vector
Definition 5 (Moreau enhanced model [29, 30])
Given a convex function $f\in\Gamma_{0}({\mathcal{H}})$, its Moreau enhanced model of index $\lambda\in{\mathbb{R}}_{++}$ is defined by
$f_{\lambda}:{\mathcal{H}}\rightarrow(-\infty,+\infty]:x\mapsto f(x)-~{}^{\lambda}f(x).$ (41)
[…]
Example 1 (Firm-shrinkage for vector)
[…]
$T^{\rm v\text{-}firm}_{\lambda_{1},\lambda_{2}}:{\mathcal{H}}\rightarrow{\mathcal{H}}:(0\neq)x\mapsto\frac{x}{\left\|x\right\|}{\rm firm}_{\lambda_{1},\lambda_{2}}(\left\|x\right\|)$ (42)
with $T^{\rm v\text{-}firm}_{\lambda_{1},\lambda_{2}}(0):=0$. Then, $T^{\rm v\text{-}firm}_{\lambda_{1},\lambda_{2}}$ is the proximity operator of the ($\lambda_{1}/\lambda_{2}$)-weakly convex function $\lambda_{1}(\left\|\cdot\right\|)_{\lambda_{2}}$, where $(\left\|\cdot\right\|)_{\lambda_{2}}(:=\left\|\cdot\right\|-~{}^{\lambda_{2}}\!\left\|\cdot\right\|)=\varphi_{\lambda_{2}}^{\rm MC}\circ\left\|\cdot\right\|$.
Example 2 (Group firm-shrinkage)
Let ${\mathcal{H}}_{1}$, ${\mathcal{H}}_{2}$, $\cdots$, ${\mathcal{H}}_{G}$ be real Hilbert spaces equipped with the norms $\left\|\cdot\right\|_{{\mathcal{H}}_{1}}$, $\left\|\cdot\right\|_{{\mathcal{H}}_{2}}$, $\cdots$, $\left\|\cdot\right\|_{{\mathcal{H}}_{G}}$, respectively, and define the product space ${\mathcal{H}}:={\mathcal{H}}_{1}\times{\mathcal{H}}_{2}\times\cdots\times{\mathcal{H}}_{G}$ equipped with the norm $\left\|x\right\|_{\mathcal{H}}:=\sqrt{\sum_{i=1}^{G}\left\|x_{i}\right\|_{{\mathcal{H}}_{i}}^{2}}$, $x:=(x_{i})_{i=1}^{G}\in{\mathcal{H}}$. Given constants $\lambda_{1}\in{\mathbb{R}}_{++}$ and $\lambda_{2}\in(\lambda_{1},+\infty)$, we define the group firm-shrinkage operator:
$T^{\rm g\text{-}firm}_{\lambda_{1},\lambda_{2}}:{\mathcal{H}}\rightarrow{\mathcal{H}}:(x_{i})_{i=1}^{G}\mapsto\left(T^{\rm v\text{-}firm}_{\lambda_{1},\lambda_{2}}(x_{i})\right)_{i=1}^{G},$ (43)
where $T^{\rm v\text{-}firm}_{\lambda_{1},\lambda_{2}}$ is defined by (42) in each space. Then, $T^{\rm g\text{-}firm}_{\lambda_{1},\lambda_{2}}$ is the proximity operator of the $(\lambda_{1}/\lambda_{2})$-weakly convex function $\lambda_{1}(\left\|\cdot\right\|_{{\rm g},1})_{\lambda_{2}}$, where $\left\|\cdot\right\|_{{\rm g},1}:{\mathcal{H}}\rightarrow{\mathbb{R}}_{+}:(x_{i})_{i=1}^{G}\mapsto\sum_{i=1}^{G}\left\|x_{i}\right\|_{{\mathcal{H}}_{i}}$.
The proofs of Examples 1 and 2 are given in Appendices C and D, respectively.
Example 3 (Neural Network with Tied Weights)
Let $\nabla\psi:{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{m}$ be a firmly-nonexpansive activation operator (see Section II-C). In fact, many of the known activation functions are firmly nonexpansive [2], such as ReLU, sigmoid, and softmax. For instance, ReLU ${\mathbb{R}}\rightarrow{\mathbb{R}}:x\mapsto\begin{cases}x&\mbox{ if }x\in{\mathbb{R}}_{+}\\ 0&\mbox{ otherwise}\end{cases}$ can be expressed as the proximity operator ${\rm s\mbox{-}Prox}_{\varphi}$ of the indicator function $\varphi:=\iota_{{\mathbb{R}}_{+}}:{\mathbb{R}}\rightarrow\{0,+\infty\}:x\mapsto\begin{cases}0&\mbox{ if }x\in{\mathbb{R}}_{+}\\ +\infty&\mbox{ otherwise,}\end{cases}$ where ${\rm s\mbox{-}Prox}_{\varphi}$ is firmly nonexpansive owing to the convexity of $\iota_{{\mathbb{R}}_{+}}$ (see Section II-E). Let ${\boldsymbol{W}}\in{\mathbb{R}}^{m\times n}$ be a nonzero weight matrix. Then, a weight-tied neural network [56] $T:={\boldsymbol{W}}^{\top}\circ\nabla\psi\circ{\boldsymbol{W}}=\nabla(\psi\circ{\boldsymbol{W}})$ can be expressed as $T={\rm s\mbox{-}Prox}_{\varphi}$ with $\varphi:=(\psi\circ{\boldsymbol{W}})^{*}-(1/2)\left\|\cdot\right\|_{2}^{2}$. Here, if $\kappa:=\|{\boldsymbol{W}}^{\top}{\boldsymbol{W}}\|>1$, $T$ is a MoL-Grad denoiser, and $\varphi$ is $(1-\kappa^{-1})$-weakly convex. If $\kappa\in(0,1]$, in contrast, $T$ is Moreau’s proximity operator of the convex function $\varphi$. More practical network architectures are discussed in [57], where multi-layered neural networks with tied nonnegative weights involving residual skip connections are shown to be MoL-Grad denoisers.
Example 4 (Debiased OSCAR)
Given constants $\lambda_{1},\lambda_{2}\in{\mathbb{R}}_{++}$, the octagonal shrinkage and clustering algorithm for regression (OSCAR) [58, 59, 60] is defined by
$\Omega_{\lambda_{1},\lambda_{2}}^{{\rm OSCAR}}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}:{\boldsymbol{x}}\mapsto\lambda_{1}\left\|{\boldsymbol{x}}\right\|_{1}+\lambda_{2}\sum_{i<j}\max\{\left|x_{i}\right|,\left|x_{j}\right|\}.$
For $\omega,\eta\in(1,+\infty)$, the operator
$T_{\lambda_{1},\lambda_{2},\omega,\eta}^{{\rm DOSCAR}}:=\omega{\rm s\mbox{-}Prox}_{\Omega_{\lambda_{1},\lambda_{2}}^{{\rm OSCAR}}}-(1-\omega){\rm s\mbox{-}Prox}_{\eta\Omega_{\lambda_{1},\lambda_{2}}^{{\rm OSCAR}}}$ (44)
is called the debiased OSCAR (DOSCAR) [51]. The operator $T_{\lambda_{1},\lambda_{2},\omega,\eta}^{{\rm DOSCAR}}$ is the proximity operator of a certain weakly convex function.
Finally, we mention that a principled way of deriving a continuous relaxation of a given discontinuous shrinkage operator is discussed under the MoL-Grad framework in [61], giving another example of the MoL-Grad denoiser.
V-C Prior Art I: Plug-and-Play Methods
Convergence analysis of plug-and-play (PnP) method: There exist a significant amount of works concerning the theoretical properties of the PnP method (see [62, 63, 64] for extensive lists of references). Among many others, [65] has proved the convergence of the PnP forward-backward splitting (PnP-FBS) method and the PnP Douglas-Rachford splitting (PnP-DRS) method, where every single iterate is assumed contraction by restricting the Lipschitz constant of ${\rm Id}-T$ with the so-named real spectral normalization (realSN). Here, the Lipschitz constant must be smaller, for instance, than $\varepsilon:=2\rho/(\kappa-\rho)$ in the case of the PnP-FBS method (see Theorem 2 for definition of $\rho$ and $\kappa$). In this framework, the limit point of the iterate is the unique fixed point of the contraction mapping (Banach-Picard).
In [66], the class of averaged-nonexpansive denoiser has been studied, encompassing those denoisers which cannot be expressed in terms of the proximity operator. In this framework, the limit point is characterized as a fixed point of the iterate. In those works, the implicit regularizer (if exists) has not been given in an explicit form. There are several previous works considering the implicit regularizer: the so-called gradient-step denoiser and the regularization by denoising approach. Those two approaches will be discussed below.
TABLE I: Comparisons of the existing denoisers for the plug-and-play method in terms of (i) the condition on the Lipschitz constant L()($\cdot$)( ⋅ ), (ii) smoothness of the implicit regularizer $\varphi$, and (iii) the range of the weak-convexity constant of $\varphi$
method Lipschitz condition smoothness range of ($1-\beta$)
realSN [65] L$({\rm Id}-T)<\varepsilon$
averaged [66] L$(T)\leq 1$
GS [37] L$({\rm Id}-T)<1$ smoooth $(0,1/2)$
RED [67] L$(T)\leq 1$
MMO [35] L$(2T-{\rm Id})\leq 1$
MoL-Grad free free $(0,1)$
Gradient-step (GS) denoiser: The GS denoiser has been studied in the literature [68, 69, 37]. Among those studies, the most relevant results would be the one in [37, Proposition 3.1], where ${\rm Id}-\nabla\psi$ is assumed to be contraction ($\kappa$-Lipschitz continuous for $\kappa<1$).101010 The arguments of [37, Proposition 3.1] rely essentially on [70]. Because of this restrictive assumption, the result in [37, Proposition 3.1] has several limitations: (i) $T$ possesses a positive definite Jacobian matrix at every point, and (ii) $\varphi$ is smooth and $\eta_{\rm GS}$-weakly convex for $\eta_{\rm GS}\in(0,1/2)$ ($\psi$ is strongly convex). In particular, the smoothness of $\varphi$ could be a strict limitation (see Remark 5 below). The implicit regularizer given in [37, Proposition 3.1] can be reproduced as a special case of Theorem 1 straightforwardly.
In [71], the convergence of the PnP method with the minimum mean squared error (MMSE) denoiser has been studied, where the MMSE denoiser is expressed in a form of GS denoiser through Tweedie’s formula. The implicit regularizer is thus given in the same way as in [37, Proposition 3.1], and it must be smooth accordingly.
Regularization by denoising (RED): The RED approach [67] relates specific algorithms to a variational optimization problem. We emphasize that it is not such an approach that relates a denoiser itself directly to an implicit regularizer. In this approach, the denoiser $T:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ is assumed to possess (i) nonexpansiveness, (ii) symmetric Jacobian $JT({\boldsymbol{x}})$ at every point ${\boldsymbol{x}}\in{\mathcal{H}}$, and (iii) the so-called local homogeneity: $T(c{\boldsymbol{x}})=cT({\boldsymbol{x}})$ for every $c\in[1-\epsilon,1+\epsilon]$ for a sufficiently small $\epsilon>0$. Here, (iii) is a key assumption to ensure $T({\boldsymbol{x}})=[JT({\boldsymbol{x}})]{\boldsymbol{x}}$, from which it can be shown that $\nabla h_{\rm RED}({\boldsymbol{x}})={\boldsymbol{x}}-T({\boldsymbol{x}})$ for $h_{\rm RED}({\boldsymbol{x}}):=(1/2){\boldsymbol{x}}^{\top}({\boldsymbol{x}}-T({\boldsymbol{x}}))$. We stress here that the relation between the denoiser $T$ and the regularizer $h_{\rm RED}$ is essentially different from our case. It has been pointed out that the RED algorithm does not actually minimize the variational objective [72], and another approach called consensus equilibrium has also been proposed as an optimization-free generalization [73].
Characterization of the limit point for plug-and-play method: In [35, Proposition 3.1], a characterization of the limit point of the vector sequence generated by the forward-backward splitting algorithm using the plug-and-play method has been presented through variational inequalities for a general family of maximally monotone operator. While the scope therein is wider in the sense that their analysis covers a general case rather than focusing solely on optimization, the operator considered therein is neural network denoisers. In addition, we explicitly showed in Section IV-B that, when the plug-and-play is applied to the primal-dual splitting algorithm, it is still possible to ensure the convergence to an optimal point with a slight modification. We repeat that the operator is not necessarily nonexpansive in our case, whereas it is supposed to be firmly nonexpansive in [35, Proposition 3.1].
Remark 5 (Novelty of MoL-Grad denoiser)
Table I shows comparisons of the proposed and existing approaches. To the best of the authors’ knowledge, MoL-Grad is the first denoiser that requires no control of its Lipschitz constant to guarantee the convergence of the operator-regularization (PnP) algorithms. In addition, the implicit regularizer $\varphi(=\psi^{*}-(1/2)\left\|\cdot\right\|^{2})$ induced by MoL-Grad denoisers is not limited to smooth ones. This means that the class of MoL-Grad denoiser is sufficiently large to contain those denoisers inducing nonsmooth regularizers. Note here that, for $\varsigma\in{\mathbb{R}}_{++}$, (i) $\varphi$ is $\varsigma$-smooth $\Rightarrow$ (ii) $\varphi+(1/2)\left\|\cdot\right\|^{2}(=\psi^{*})$ is $(\varsigma+1)$-smooth $\Leftrightarrow$ (iii) $\psi$ is $1/(\varsigma+1)$-strongly convex $\Rightarrow$ (iv) $T=\nabla\psi$ is $1/(\varsigma+1)$-strongly monotone (see Section II-B for the equivalence (ii) $\Leftrightarrow$ (iii)). This implies that the MoL-Grad denoiser $T$ needs not to be strongly monotone. Nonsmoothness of regularizers has actually played a crucial role in sparse modeling. For instance, the firm/garrote shrinkage presented in Section V-A induces the nonsmooth regularizer given in (38) or (40). For the weight-tied neural networks discussed in Example 3, moreover, adoption of a saturating activation function in the network such as ReLU makes the associated regularizer nonsmooth, because $\nabla^{2}\psi({\boldsymbol{y}})=JT({\boldsymbol{y}})={\boldsymbol{O}}$ at some ${\boldsymbol{y}}\in{\mathbb{R}}^{m}$.
V-D Prior Art II: Other Related Works
Characterization of implicit regularizer: In [2], the softmax activation function is shown to be Moreau’s proximity operator of a certain convex function by using [39, Corollary 24.5], which is related to, but is different from, Theorem 1(b). Specifically, [39, Corollary 24.5] states that, given $\varphi\in\Gamma_{0}({\mathcal{H}})$ and $f\in\Gamma_{0}({\mathcal{H}})$ such that $\varphi=f-(1/2)\left\|\cdot\right\|^{2}$, it holds that ${\rm s\mbox{-}Prox}_{\varphi}=\nabla f^{*}$. Here, $\varphi$ is implicitly assumed to be $1$-strongly convex. Meanwhile, by changing $\psi^{*}$ in Theorem 1(b) by $f$, it can essentially be stated as follows: given $\varphi+((1-\beta)/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ and $f-(\beta/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ with $\varphi=f-(1/2)\left\|\cdot\right\|^{2}$, it holds that ${\rm s\mbox{-}Prox}_{\varphi}=\nabla f^{*}$. Here, $\varphi$ is $(1-\beta)$-weakly convex (while $f$ is $\beta$-strongly convex) in contrast to the former case.
Proximity operator of nonconvex function: In [46], it has been shown that, given a proper lower semicontinuous function, it is weakly convex if and only if its proximity operator is cocoercive. The arguments therein are based on the notion of abstract subgradient, while our arguments are based solely on the standard subgradient (see (2)) adopted in convex analysis. In [45, Proposition 2], given a mapping $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ and $L>0$, the equivalence of the following two statements has been established: (a) $T(x)\in\mathbf{Prox}_{\varphi}(x)$, $\forall x\in{\mathcal{H}}$, for some $\varphi:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ such that $\varphi+(1/2)[1-1/L]\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ (see Appendix A for the definition of set-valued operator $\mathbf{Prox}_{\varphi}$), and (b) $T$ is $L$-Lipschitz continuous and $T(x)\in\partial\psi(x)$, $\forall x\in{\mathcal{H}}$, for some $\psi\in\Gamma_{0}({\mathcal{H}})$. In [45], the proximity operator is defined as a set-valued operator (or its selection), and the optimization aspect has not been discussed.
In contrast to the studies in [46, 45], our s-prox operator is defined as a “unique” minimizer of the penalized function (see Definition 4). This is because our primal focus is on the explainability perspective of optimization algorithms, and because for this reason our denoiser is continuous (see Proposition 1). Theorem 1 presented in Section III-A is a refinement of [45, Proposition 2]. Specifically, Theorem 1 explicitly shows that $\mathbf{Prox}_{\varphi}$ is actually single-valued, and it discloses the exact relation between $\varphi$ and $\psi$, based on which the characterization of implicit regularizers is given for the operator-regularization approaches in Section IV.
Proximity operator of weakly convex function in a linear inverse problem: In [52], the proximity operator (in the sense of Definition 4 essentially) of weakly convex functions has been studied in a linear inverse problem, and the convergence of the forward-backward splitting algorithm with the s-prox operator has been analyzed. While the main results therein are related to Theorem 2, the present study has a wider scope. Specifically, it includes Theorem 1, Propositions 2 and 3, and Theorem 3 as well as the discussions in Section IV-C to make the idea be applicable to other operator splitting algorithms. Moreover, the study in [52] considers the case in which (i) ${\mathcal{H}}$ is the Euclidean space, and (ii) the range of $\varphi$ is ${\mathbb{R}}$. whereas the present study concerns the case in which (i) ${\mathcal{H}}$ is a general Hilbert space, and (ii) $\varphi$ may take the value +$\infty$+ ∞.
Other works: There are some other related works. In [33], a class of “separable” shrinkage operators (i.e., one-dimensional shrinkage operators essentially) and their induced penalty functions have been studied. In [74], a study on one-dimensional monotone operators has been presented, showing that “a non-decreasing non-constant function with at most a countable number of discontinuities” can be expressed as a “selection” of the (set-valued) proximity operator of a weakly convex function. Although those studies are related to Theorem 1, the scope of the present study is different in the sense that it concerns the operator-regularization approaches explicitly based on the s-prox operator of nonseparable weakly-convex function defined on a Hilbert space.
VI Simulations
Simulations are conducted to show (i) how the theory works in practice and (ii) how the generalized proximity operator competes with Moreau’s one. For this purpose, the firm shrinkage operator will be adopted as the MoL-Grad denoiser $T$ in Algorithm 1.
VI-A Firm Shrinkage Plugged into Algorithm 1
We consider the linear system model ${\boldsymbol{y}}:={\boldsymbol{A}}{\boldsymbol{x}}_{\diamond}+{\boldsymbol{\varepsilon}}\in{\mathbb{R}}^{m}$, where ${\boldsymbol{x}}_{\diamond}\in{\mathcal{H}}:={\mathbb{R}}^{n}$ is the $n$-dimensional Euclidean vector to estimate, ${\boldsymbol{A}}\in{\mathbb{R}}^{m\times n}$ is the sensing matrix, and ${\boldsymbol{\varepsilon}}\in{\mathbb{R}}^{m}$ is the noise vector. We consider the overdetermined case $\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})>0$, where $\lambda_{\min}(\cdot)$ denotes the smallest eigenvalue. (See [51] for the case of $\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})=0$.) We suppose that ${\boldsymbol{x}}_{\diamond}$ is piecewise constant so that the difference vector ${\boldsymbol{D}}{\boldsymbol{x}}=[x_{1}-x_{2},x_{2}-x_{3},\cdots,x_{n-1}-x_{n}]^{\top}$ is sparse, where ${\boldsymbol{D}}:=[{\boldsymbol{I}}~{}{\boldsymbol{0}}]-[{\boldsymbol{0}}~{}{\boldsymbol{I}}]\in{\mathbb{R}}^{(n-1)\times n}$ with the identity matrix ${\boldsymbol{I}}$ and the zero vector ${\boldsymbol{0}}$ of length $n-1$. We therefore let $L:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n-1}=:\mathcal{U}:{\boldsymbol{x}}\mapsto{\boldsymbol{D}}{\boldsymbol{x}}$. The data fidelity term is set to $f({\boldsymbol{x}}):=(1/2)\left\|{\boldsymbol{A}}{\boldsymbol{x}}-{\boldsymbol{y}}\right\|^{2}_{2}$. In this case, since the Hessian matrices of $f$ and $\hat{f}$ are given by $\nabla^{2}f({\boldsymbol{x}})={\boldsymbol{A}}^{\top}{\boldsymbol{A}}$ and $\nabla^{2}\hat{f}({\boldsymbol{x}})={\boldsymbol{A}}^{\top}{\boldsymbol{A}}-(\rho/\left\|{\boldsymbol{D}}\right\|^{2}){\boldsymbol{D}}^{\top}{\boldsymbol{D}}$, respectively, at every ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, $f$ is $\rho$-strongly convex for $\rho:=\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})$, and $\hat{f}$ is $\kappa$-smooth for $\kappa:=\|{\boldsymbol{A}}^{\top}{\boldsymbol{A}}-(\rho/\left\|{\boldsymbol{D}}\right\|^{2}){\boldsymbol{D}}^{\top}{\boldsymbol{D}}\|$.
Recall that the firm shrinkage given in (37) has two parameters $\lambda_{1}$ and $\lambda_{2}$, and its corresponding weakly convex function is $\varphi:=\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}$ with $\varphi+(\lambda_{1}/(2\lambda_{2}))\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathbb{R}}^{n})$, which means that $\beta=1-\lambda_{1}/\lambda_{2}$. Hence, in view of Theorem 3, under an appropriate setting, the sequence $({\boldsymbol{x}}_{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 converges to a solution of the following minimization problem:
$\min_{{\boldsymbol{x}}\in{\mathbb{R}}^{n}}f({\boldsymbol{x}})+[\sigma+\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})/\left\|{\boldsymbol{D}}\right\|^{2}]\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}({\boldsymbol{D}}{\boldsymbol{x}}).$ (45)
From (45), the impact of the regularizer $\varphi_{\lambda_{2}}^{\rm MC}$ scales with $\lambda_{1}$ as well as the dual step size $\sigma$.
As mentioned in Remark 4, the $\sigma$ satisfying condition (i) in Theorem 3 can be parametrized by $\delta\in(0,1]$ as $\sigma:=\delta(\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})/\left\|{\boldsymbol{D}}\right\|^{2})[(\lambda_{2}/\lambda_{1})-1]>0$, making (45) into
$\min_{{\boldsymbol{x}}\in{\mathbb{R}}^{n}}f({\boldsymbol{x}})+\mu^{-1}\varphi_{\lambda_{2}}^{\rm MC}({\boldsymbol{D}}{\boldsymbol{x}}),$ (46)
where $\mu:=\left\|{\boldsymbol{D}}\right\|^{2}/[\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})((1-\delta)\lambda_{1}+\delta\lambda_{2})]>0$. Here, $\left\|{\boldsymbol{D}}\right\|^{2}$ and $\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})$ serve as a sort of normalization, while $(1-\delta)\lambda_{1}+\delta\lambda_{2}\in(\lambda_{1},\lambda_{2}]$ weights the regularizer $\varphi_{\lambda_{2}}^{\rm MC}$ relative to $f$; increasing $\delta$ shifts the weight from $\lambda_{1}$ to $\lambda_{2}$.
Letting $\delta:=1$, or equivalently setting $\sigma$ to its upper bound, reduces (45) to
$\min_{{\boldsymbol{x}}\in{\mathbb{R}}^{n}}f({\boldsymbol{x}})+\frac{\lambda_{2}\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})}{\left\|{\boldsymbol{D}}\right\|^{2}}\varphi_{\lambda_{2}}^{\rm MC}({\boldsymbol{D}}{\boldsymbol{x}}),$ (47)
which is independent of $\lambda_{1}$. Note here that the dependence on $\lambda_{1}$ is eliminated by condition (i) to guarantee convexity of the overall cost $f+g\circ L$. As such, the performance of Algorithm 1 in this case is governed solely by $\lambda_{2}$, given the matrices ${\boldsymbol{A}}$ and ${\boldsymbol{D}}$. This might be advantageous in practice, because $\lambda_{2}$ is the only parameter to tune, unless one cares the speed of convergence.
VI-B Experimental Verification of Theorem 3
Since $\varphi_{\lambda_{2}}^{\rm MC}=\left\|\cdot\right\|_{1}-~{}^{\lambda_{2}}(\left\|\cdot\right\|_{1})$, (46) can be rewritten as follows:
$\min_{{\boldsymbol{x}}\in{\mathbb{R}}^{n}}\underbrace{\mu f({\boldsymbol{x}})-~{}^{\lambda_{2}}(\left\|\cdot\right\|_{1})({\boldsymbol{D}}{\boldsymbol{x}})}_{=\tilde{f}({\boldsymbol{x}})}+\left\|{\boldsymbol{D}}{\boldsymbol{x}}\right\|_{1}.$ (48)
As $\tilde{f}:=\mu f-~{}^{\lambda_{2}}(\left\|\cdot\right\|_{1})\circ{\boldsymbol{D}}$ is a smooth convex function with its gradient available in a closed form, the minimization problem in (48) can be solved by the operator splitting algorithm such as the primal-dual splitting method with convergence guarantee.111111 This is possible in this specific example, because the function $\varphi$ corresponding to the firm shrinkage is available. In general, $\varphi$ could be unavailable, and thus this kind of approach cannot be used in such a case. To demonstrate the validity of Theorem 3 by simulation, we apply the Condat–Vũ algorithm (form II) to (48) to see whether Algorithm 1 converges to the same point. We show, more specifically, that the discrepancy is vanishing between $({\boldsymbol{x}}_{k},{\boldsymbol{u}}_{k})$ of Algorithm 1 and $(\tilde{{\boldsymbol{x}}}_{k},\tilde{{\boldsymbol{u}}}_{k})$ of the Condat–Vũ algorithm applied to (48), where “discrepancy” is quantified as follows:
$\frac{\left\|{\boldsymbol{x}}_{k}-\tilde{{\boldsymbol{x}}}_{k}\right\|_{2}^{2}+\left\|{\boldsymbol{u}}_{k}-\tilde{{\boldsymbol{u}}}_{k}\right\|_{2}^{2}}{\left\|{\boldsymbol{x}}_{k}\right\|_{2}^{2}+\left\|{\boldsymbol{u}}_{k}\right\|_{2}^{2}}.$ (49)
We let $n:=256$, $m=1024$, and ${\boldsymbol{x}}_{\diamond}$ depicted in Fig. 3 is used. For Algorithm 1, we set $\delta:=1$, $\tau:=0.9(\sigma\left\|{\boldsymbol{D}}\right\|^{2}+\kappa/2)^{-1}$, $\lambda_{1}:=2.5$, and $\lambda_{2}:=5$. We repeat that $\lambda_{1}$ does not change the solution, see (47). For the approach based on (48), we set $\sigma:=0.2$, and all other parameters are the same as for Algorithm 1. Figure 4 plots the discrepancy given in (49) (which is labeled as ‘${\boldsymbol{x}},{\boldsymbol{u}}$’); the discrepancy in terms of the primal variables and that in terms of the dual variables are also plotted, labeled as ‘${\boldsymbol{x}}$’ and ‘${\boldsymbol{u}}$’, respectively. It can be seen that the discrepancy vanishes as time goes by.
For a comparison, we also test a straightforward approach (which has no theoretical guarantee): apply the Condat-Vũ algorithm (form II) directly to the nonconvex formulation (46), which is essentially the same as (48). Here, ${\rm s\mbox{-}Prox}_{\sigma g^{*}}$ appearing in the algorithm is replaced121212 This replacement is heuristic, and there is no theoretical support for that. by ${\rm Id}-\sigma{\rm s\mbox{-}Prox}_{\varphi_{\lambda_{2}}^{\rm MC}/(\mu\sigma)}\circ(\sigma^{-1}{\rm Id})$, where ${\rm s\mbox{-}Prox}_{\varphi_{\lambda_{2}}^{\rm MC}/(\mu\sigma)}$ is the proximity operator in the sense of Definition 4, which coincides with the firm shrinkage with $\lambda_{1}:=1/(\mu\sigma)$. The $\sigma$ here needs to be chosen so that $\lambda_{1}<\lambda_{2}$. The results are plotted in Fig. 5. It can be seen that the discrepancy (‘${\boldsymbol{x}},{\boldsymbol{u}}$’) does not vanish. Although the discrepancy in terms of ‘${\boldsymbol{x}}$’ vanishes, there is no theoretical guarantee for that, to the best of authors’ knowledge.
VI-C Performance Comparisons
We compare the performance of the approach described in Section VI-A to that of the $\ell_{1}$-based total variation approach under the same setting basically as in Section VI-B. For the former approach, since $\lambda_{2}$ is the only parameter that affects the solution, we change $\lambda_{2}$ to see how the performance changes accordingly. The cost function of the latter approach is given by $\mu f+\left\|\cdot\right\|_{1}\circ{\boldsymbol{D}}$, and we change $\mu$ for this approach. The evaluation metric is system mismatch defined by $\left\|{\boldsymbol{x}}_{\diamond}-{\boldsymbol{x}}_{k}\right\|_{2}^{2}/\left\|{\boldsymbol{x}}_{\diamond}\right\|_{2}^{2}$.
Figure 6 plots the results averaged over 300 independent trials. The approach described in Section VI-A outperforms the $\ell_{1}$-based approach. Note here that it is a fair comparison in the sense that both approaches have a single tuning parameter. To see the full potential of the firm shrinkage operator, one may also tune $\lambda_{1}$ and $\delta$, and this will lead to further performance improvements at the cost of additional parameter tuning.
[…]
Figure 3: The vector ${\boldsymbol{x}}_{\diamond}$ used in the simulation.
[…]
Figure 4: Agreement of the learning curves in line with Theorem 3.
[…]
Figure 5: Disagreement of the learning curves for the straightforward approach.
[…]
Figure 6: Performance comparisons: convex (soft shrinkage) and weakly-convex (firm shrinkage).
VII Concluding Remarks
This paper studied the operator-regularization approaches from the “optimality” perspective based on the “single-valued” proximity operator of nonconvex functions. The single-valuedness here was a necessary condition for the operator to be continuous. (Proposition 1). The fundamental result was presented, showing that a given operator is the “single-valued” proximity operator of a weakly convex function if and only if it is a MoL-Grad denoiser, i.e., the gradient of a $\kappa$-smooth convex function for $\kappa>1$ (Theorem 1). Extensions of Moreau’s decomposition were also presented, expressing a vector in terms of the proximity operators of a weakly convex function and the conjugate of its convexified one (Propositions 2 and 3). In addition to the forward-backward splitting algorithm, the primal-dual splitting algorithm employing a MoL-Grad denoiser (under the slight modification) was shown to generate a vector sequence convergent weakly to a minimizer of the cost function involving the implicit regularizer (Theorems 2 and 3). The simulation results matched our theoretical findings as well as demonstrating potential advantages of the proposed approach.
The proposed framework has two remarkable advantages over the previous works on convergence analysis of the PnP method. First, it is free from Lipschitz-constant control during the learning of the denoiser. Second, the implicit regularizer induced by the MoL-Grad denoiser is not confined to smooth ones, meaning that our framework encompasses a large class of denoisers inducing possibly nonsmooth regularizers. An important implication of our findings is the following: when the fidelity term is strongly convex and when such a modified operator splitting algorithm that has guarantee of convergence to an optimal point is desired, one can restrict attention to MoL-Grad denoisers, or, equivalently, (the proximity operator of) weakly convex functions. Applications of the presented theory to various problems are left as an exciting open issue.
Appendix A Proof of Proposition 1
We first show the monotonicity of the “set-valued” operator $\mathbf{Prox}_{f}:{\mathcal{H}}\rightarrow 2^{\mathcal{H}}:x\mapsto\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}J_{x}(y)$, where $J_{x}:=f+(1/2)\left\|x-\cdot\right\|^{2}$. By virtue of Fermat’s rule together with Lemma 1 given in Section III-A, we have
$p\!\in\!\mathbf{Prox}_{f}(x)\Leftrightarrow 0\!\in\!\partial J_{x}(p)\Leftrightarrow x\!\in\!\partial J_{0}(p)\Leftrightarrow p\!\in\!\big{(}\partial J_{0}\big{)}^{-1}(x),$
which implies that $\mathbf{Prox}_{f}=\big{(}\partial J_{0}\big{)}^{-1}$. As the subdifferential $\partial J_{0}$ is monotone [39, Example 20.3], its inverse $\big{(}\partial J_{0}\big{)}^{-1}$ is also monotone [39, Proposition 20.10].
For contradiction, suppose, for an arbitrarily fixed $x\in{\mathcal{H}}$, that $\mathbf{Prox}_{f}(x)$ contains two distinct vectors $T(x)$ and $T(x)+\delta$ for some $\delta\in{\mathcal{H}}\setminus\{0\}$. The monotonicity of $\mathbf{Prox}_{f}$ suggests that, for every ϵ++italic-$\epsilon\in{\mathbb{R}}_{++}$, it holds that
$~{}\left\langle{T(x+\epsilon\delta)-(T(x)+\delta)},{(x+\epsilon\delta)-x}\right\rangle\geq 0$
$\Leftrightarrow$ $~{}\left\langle{T(x+\epsilon\delta)-T(x)-\delta},{\epsilon\delta}\right\rangle\geq 0$
$\Leftrightarrow$ $~{}\left\langle{T(x+\epsilon\delta)-T(x)},{\delta}\right\rangle\geq\left\|\delta\right\|^{2}.$ (A.1)
By the continuity of $T$ as well as that of the inner product, (A.1) reads
$0=\lim_{\epsilon\downarrow 0}\left\langle{T(x+\epsilon\delta)-T(x)},{\delta}\right\rangle\geq\left\|\delta\right\|^{2}>0,$ (A.2)
which gives contradiction. Hence, $\mathbf{Prox}_{f}$ is a single-valued operator over ${\mathcal{H}}$, meaning that $J_{x}$ has a unique minimizer for every $x\in{\mathcal{H}}$. $\blacksquare$
Appendix B Preservation of averaged nonexpansiveness
Lemma B.1
Let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$. Then, for every $\alpha\in(0,1)$ and ϱ++italic-$\varrho\in{\mathbb{R}}_{++}$, the following hold.
  1. 1.
    $T$ is $\alpha$-averaged nonexpansive if and only if $(c^{-1}T)\circ(c{\rm Id})$ is $\alpha$-averaged nonexpansive for every $c\in{\mathbb{R}}_{++}$.
  2. 2.
    $\varrho T$ is $\alpha$-averaged nonexpansive if and only if $T\circ(\varrho{\rm Id})$ is $\alpha$-averaged nonexpansive. In particular, $T$ is $\varrho$-cocoercive if and only if $T\circ(\varrho{\rm Id})$ is firmly ((1/2)-averaged) nonexpansive.
Proof.
1): It can be verified that
$T=(1-\alpha){\rm Id}+\alpha N,~{}\exists N:{\mathcal{H}}\rightarrow{\mathcal{H}}\mbox{ nonexpansive}$
$\Leftrightarrow$ $~{}c^{-1}T\circ(c{\rm Id})=(1-\alpha){\rm Id}+\alpha c^{-1}N\circ(c{\rm Id}),$
$\hskip 80.00012pt\exists N:{\mathcal{H}}\rightarrow{\mathcal{H}}\mbox{ nonexpansive.}$. (B.3)
Here, $\tilde{N}:=c^{-1}N\circ(c{\rm Id})$ is nonexpansive if $N$ is nonexpansive, because $\left\|\tilde{N}(x)-\tilde{N}(y)\right\|=c^{-1}\left\|N(cx)-N(cy)\right\|\leq c^{-1}\left\|cx-cy\right\|=\left\|x-y\right\|$. The converse can be verified in an analogous way. Hence, (B.3) implies the equivalence of the $\alpha$-averaged nonexpansiveness of $T$ and $c^{-1}T\circ(c{\rm Id})$ for an arbitrary $c>0$.
2): As the second claim can be immediately verified by letting $\alpha:=1/2$ in the first one, we prove the first claim in the following. Assume that $\varrho T$ is $\alpha$-averaged nonexpansive. Then, by Lemma B.1.1, $c^{-1}\varrho T\circ(c{\rm Id})$ is $\alpha$-averaged nonexpansive for every $c\in{\mathbb{R}}_{++}$. The specific choice of $c:=\varrho$ yields $\alpha$-averaged nonexpansiveness of $T\circ(\varrho{\rm Id})$.
To verify the reverse implication, assume that $T\circ(\varrho{\rm Id})$ is $\alpha$-averaged nonexpansive. Then, by Lemma B.1.1 again, $c^{-1}T\circ(c\varrho{\rm Id})$ is $\alpha$-averaged nonexpansive for every $c\in{\mathbb{R}}_{++}$. The specific choice of $c:=\varrho^{-1}$ yields $\alpha$-averaged nonexpansiveness of $\varrho T$. ∎
Appendix C Proof of Example 1
We first prove the following lemma.
Lemma C.1
[…]
Proof of Lemma C.1: By definition, we have
(${}^{\lambda}(h\circ\left\|\cdot\right\|)(x)=\min_{y\in{\mathcal{H}}}\Big{(}h(\left\|y\right\|)+\frac{1}{2\lambda}\left\|y-x\right\|^{2}\Big{)}.$ (C.4)
If $x=0$, (${}^{\lambda}(h\circ\left\|\cdot\right\|)(0)=~{}^{\lambda}(h\circ\left|\cdot\right|)(\left\|0\right\|)=h(0)$. We now assume that $x\neq 0$. Based on the orthogonal decomposition of ${\mathcal{H}}=\mathcal{M}\oplus\mathcal{M}^{\perp}$ with $\mathcal{M}:={\rm span}\{x\}$ and its orthogonal complement $\mathcal{M}^{\perp}$, every $y$ can be represented by $y=\xi x+y_{\perp}$ with $\xi\in{\mathbb{R}}$ and $y_{\perp}\in\mathcal{M}^{\perp}$. Thus, by the Pythagorean theorem, the Moreau envelope in (C.4) can be represented as follows:
($~{}^{\lambda}(h\circ\left\|\cdot\right\|)(x)=$ $~{}\min_{\xi\in{\mathbb{R}},y_{\perp}\in\mathcal{M}^{\perp}}h\Big{(}\sqrt{\left\|\xi x\right\|^{2}+\left\|y_{\perp}\right\|^{2}}\Big{)}$
$\hskip 37.00002pt+\frac{1}{2\lambda}\left(\left\|(\xi-1)x\right\|^{2}+\left\|y_{\perp}\right\|^{2}\right)$
$=$ $~{}\min_{\xi\in{\mathbb{R}}}h(\left\|\xi x\right\|)+\frac{1}{2\lambda}\left\|(\xi-1)x\right\|^{2}$
$=$ $~{}\min_{\xi\in{\mathbb{R}}}h(\left|\xi\left\|x\right\|\right|)+\frac{1}{2\lambda}(\xi\left\|x\right\|-\left\|x\right\|)^{2}$
$=$ $~{}\min_{\zeta\in{\mathbb{R}}}h(\left|\zeta\right|)+\frac{1}{2\lambda}(\zeta-\left\|x\right\|)^{2}$
$=$ ($~{}^{\lambda}(h\circ\left|\cdot\right|)(\left\|x\right\|),$ (C.5)
which proves the first claim of the lemma. Here, the second equality of (C.5) holds because the minimum is achieved only when $y_{\perp}=0$ owing to the nonincreasingness of $h$. The second claim of the lemma can be verified in an analogous way. $\blacksquare$
Proof of Example 1: Letting $h:x\mapsto x$ and $\lambda:=\lambda_{2}$ in Lemma C.1 gives ${}^{\lambda_{2}}\left\|\cdot\right\|(x)=~{}^{\lambda_{2}}\left|\cdot\right|(\left\|x\right\|)$. This, together with $(\left|\cdot\right|)_{\lambda_{2}}=\varphi_{\lambda_{2}}^{\rm MC}$, verifies that $(\left\|\cdot\right\|)_{\lambda_{2}}=\varphi_{\lambda_{2}}^{\rm MC}\circ\left\|\cdot\right\|$. Recall now that the MC function given in (38) is a symmetric nondecreasing function. Hence, letting $h=\varphi_{\lambda_{2}}^{\rm MC}$ and $\lambda:=\lambda_{1}$ in Lemma C.1 yields ${\rm s\mbox{-}Prox}_{\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}\circ\left\|\cdot\right\|}(x)=(x/\left\|x\right\|){\rm s\mbox{-}Prox}_{\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}\circ\left|\cdot\right|}(\left\|x\right\|)=(x/\left\|x\right\|){\rm s\mbox{-}Prox}_{\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}}(\left\|x\right\|)=(x/\left\|x\right\|){\rm firm}_{\lambda_{1},\lambda_{2}}(\left\|x\right\|)$, where the second equality is owing to the symmetry of $\varphi_{\lambda_{2}}^{\rm MC}$. See Section V-A and also [17, 74, 29] for the third equality. $\blacksquare$
Appendix D Proof of Example 2
Owing to the group-wise separability of the group $\ell_{1}$ norm $\left\|\cdot\right\|_{{\rm g},1}$, the Moreau envelope can be expressed in the following separable form: ${}^{\lambda_{2}}\left\|\cdot\right\|_{{\rm g},1}(x)=\min_{y\in{\mathcal{H}}}\big{[}\big{(}\sum_{i=1}^{G}\left\|y_{i}\right\|_{{\mathcal{H}}_{i}}\big{)}+(1/(2\lambda_{2}))\left\|x-y\right\|_{{\mathcal{H}}}^{2}\big{]}=\sum_{i=1}^{G}\min_{y_{i}\in{\mathcal{H}}_{i}}\big{[}\left\|y_{i}\right\|_{{\mathcal{H}}_{i}}+(1/(2\lambda_{2}))\left\|x_{i}-y_{i}\right\|_{{\mathcal{H}}_{i}}^{2}\big{]}=\sum_{i=1}^{G}{{}^{\lambda_{2}}\left\|\cdot\right\|_{{\mathcal{H}}_{i}}}(x_{i})$. Thus, the Moreau enhanced model can also be expressed in a separable form as follows: $(\left\|\cdot\right\|_{{\rm g},1})_{\lambda_{2}}(x):=\sum_{i=1}^{G}(\left\|\cdot\right\|_{{\mathcal{H}}_{i}})_{\lambda_{2}}(x_{i})$. Hence, it follows that ${\rm s\mbox{-}Prox}_{\lambda_{1}(\left\|\cdot\right\|_{{\rm g},1})_{\lambda_{2}}}(x)=\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}\sum_{i=1}^{G}[(\left\|\cdot\right\|_{{\mathcal{H}}_{i}})_{\lambda_{2}}(y_{i})+(1/(2\lambda_{1}))\left\|x_{i}-y_{i}\right\|_{{\mathcal{H}}_{i}}^{2}]=\big{(}{\rm s\mbox{-}Prox}_{\lambda_{1}(\left\|\cdot\right\|_{{\mathcal{H}}_{i}})_{\lambda_{2}}}(x_{i})\big{)}_{i=1}^{G}$, where ${\rm s\mbox{-}Prox}_{\lambda_{1}(\left\|\cdot\right\|_{{\mathcal{H}}_{i}})_{\lambda_{2}}}=T^{\rm v\text{-}firm}_{\lambda_{1},\lambda_{2}}$ by Example 1. $\blacksquare$
Appendix E Strong convexity, strong monotonicity, and cocoercivity
It is well known that the subdifferential of a strongly convex function is a strongly monotone operator. Nevertheless, the converse is not always true, and an additional condition is required for the reverse implication. On the other hand, the interplay between strong monotonicity and cocoercivity is also well understood. The following proposition summarizes those basics below. [^13] Proposition C.1 is closely related to [39, Theorem 18.15].)
Proposition C.1
Let $f\in\Gamma_{0}({\mathcal{H}})$. For every $\eta>0$, the following statements are equivalent.
  1. 1.
    $f$ is an $\eta$-strongly convex function.
  2. 2.
    $\partial f$ is $\eta$-strongly monotone operator.
  3. 3.
    $(\partial f)^{-1}:{\mathcal{H}}\rightarrow 2^{\mathcal{H}}$ is a single-valued operator with ${\rm dom}~{}(\partial f)^{-1}={\mathcal{H}}$, and $\nabla f^{*}:{\mathcal{H}}\rightarrow{\mathcal{H}}:x\mapsto u$ under $(\partial f)^{-1}(x)=\{u\}$ is an $\eta$-cocoercive operator well-defined everywhere.
  4. 4.
    $f^{*}$ is a Fréchet differentiable function on ${\mathcal{H}}$, and $\nabla f^{*}$ is $\eta^{-1}$-Lipschitz continuous with ${\rm dom}~{}\nabla f^{*}={\mathcal{H}}$.
  5. 5.
    ${\rm s\mbox{-}Prox}_{f}=(I+\partial f)^{-1}$ is a $(1+\eta)$-cocoercive operator.
Proposition C.1 will be proved by using the following three lemmas.
Lemma C.2 (Proposition 16.27 of [39])
Let $f\in\Gamma_{0}({\mathcal{H}})$. Then, ${\rm int}~{}{\rm dom}~{}f={\rm cont}~{}f\subset{\rm dom}~{}\partial f\subset{\rm dom}~{}f$.
By Lemma C.2 with $(\partial f)^{-1}=\partial f^{*}$, it holds for every $f\in\Gamma_{0}({\mathcal{H}})$ that
${\rm int}~{}{\rm dom}~{}f^{*}\subset{\rm range}~{}\partial f={\rm dom}~{}\partial f^{*}\subset{\rm dom}~{}f^{*}.$ (C.6)
Lemma C.3
Let $f\in\Gamma_{0}({\mathcal{H}})$ be strongly convex. Then, ${\rm range}~{}\partial f={\rm dom}~{}(\partial f)^{-1}={\rm dom}~{}\partial f^{*}={\mathcal{H}}$.
Proof.
The strong convexity implies the supercoercivity [^14] A function $f:{\mathcal{H}}\rightarrow[-\infty,+\infty]$ is supercoercive if $\lim_{\left\|x\right\|\rightarrow+\infty}\frac{f(x)}{\left\|x\right\|}=+\infty$; $f$ is coercive if $\lim_{\left\|x\right\|\rightarrow+\infty}f(x)=+\infty$. of $f$ by [39, Corollary 11.17], and hence ${\rm dom}~{}f^{*}={\mathcal{H}}$ by [39, Proposition 14.15]. This implies that ${\rm int}~{}{\rm dom}~{}f^{*}={\mathcal{H}}$, which together with (C.6) verifies the assertion. ∎
Lemma C.4
Let $T:{\mathcal{H}}\rightarrow 2^{\mathcal{H}}$ be a strictly monotone operator such that ${\rm range}~{}T={\mathcal{H}}$. Then, the inverse $T^{-1}$ is a single-valued operator with ${\rm dom}~{}T^{-1}={\mathcal{H}}$.
Proof.
To show that $T^{-1}$ is a single-valued operator, we suppose for contradiction that there exists a triple $(x,y,z)\in{\mathcal{H}}^{3}$ such that (i) $y\neq z$, (ii) $(x,y)\in{\rm gra}~{}T^{-1}$, and (iii) $(x,z)\in{\rm gra}~{}T^{-1}$. It then holds simultaneously that $(y,x)\in{\rm gra}~{}T$, $(z,x)\in{\rm gra}~{}T$, $y\neq z$, and $\left\langle{y-z},{x-x}\right\rangle=0$. This contradicts the definition of strictly monotone operator. It is clear that ${\rm dom}~{}T^{-1}={\rm range}~{}T={\mathcal{H}}$. ∎
Proof of Proposition C.1
1)2)1$\Rightarrow$ 2)1 ) ⇒ 2 ): It follows by [39, Example 20.25] and [39, Example 22.4].
2)3)2$\Rightarrow$ 3)2 ) ⇒ 3 ): As $\partial f$ is maximally monotone owing to $f\in\Gamma_{0}({\mathcal{H}})$, strong monotonicity of $\partial f$ implies that ${\rm range}~{}\partial f={\mathcal{H}}$ [39, Proposition 22.11]. Combining this with Lemma C.4, it can be verified that $(\partial f)^{-1}$ is a single-valued operator with ${\rm dom}~{}(\partial f)^{-1}={\mathcal{H}}$, which also means that the operator $\nabla f^{*}$ is well-defined everywhere. Note here that the strong monotonicity of $\partial f$ implies its strict monotonicity by definition. Finally, $\eta$-cocoercivity of $\nabla f^{*}$ is equivalent to $\eta$-strong monotonicity of its inverse $(\nabla f^{*})^{-1}=\partial f$ [39, Example 22.7], which completes the proof.
3)4)3$\Rightarrow$ 4)3 ) ⇒ 4 ): The $\eta$-cocoercivity of $\nabla f^{*}$ implies firm nonexpansiveness, thus 1-Lipschitz continuity, of $\eta\nabla f^{*}$, which implies again $\eta^{-1}$-Lipschitz continuity of $\nabla f^{*}$ itself. It therefore suffices to show the Fréchet differentiability of $f^{*}$. By $(\nabla f^{*})^{-1}=\partial f$ under $f\in\Gamma_{0}({\mathcal{H}})$, dom $(\partial f)^{-1}={\mathcal{H}}$ implies ${\rm dom}~{}\nabla f^{*}={\rm dom}~{}\partial f^{*}={\mathcal{H}}$, which implies again that $({\rm dom}~{}\partial f^{*}\subset){\rm dom}~{}f^{*}={\mathcal{H}}$. Thus, it follows that cont $f^{*}={\rm int}~{}{\rm dom}~{}f^{*}={\mathcal{H}}$ by Lemma C.2. This suggests that $f^{*}$ is Gateaux differentiable over ${\mathcal{H}}$ in light of [39, Proposition 17.31], which, together with the (Lipschitz) continuity of $\nabla f^{*}$, implies the Fréchet differentiability of $f^{*}$ by [39, Corollary 17.42].
4)1)4$\Rightarrow$ 1)4 ) ⇒ 1 ): Use [39, Theorem 18.15] by noting that dom $f^{*}={\mathcal{H}}$ ($\Leftarrow{\rm dom}~{}\partial f^{*}={\mathcal{H}}$).
2)5)2$\Leftrightarrow$ 5)2 ) ⇔ 5 ): Since $\partial f$ is a monotone operator, $\eta$-strong monotonicity of $\partial f$ and $(1+\eta)$-cocoercivity of $(I+\partial f)^{-1}$ are equivalent [39, Proposition 23.13]. Finally, under 5), $\partial f$ is both maximally monotone and strongly monotone, which implies ${\rm range}~{}\partial f={\mathcal{H}}$ by [39, Proposition 22.11]. $\blacksquare$
Show an operator $\partial f$ $\eta$-strongly monotone with ${\rm range}~{}\partial f\neq{\mathcal{H}}$ defined by $f\in\Gamma_{0}({\mathcal{H}})$ NOT strongly convex.
Let $f:{\mathcal{H}}\rightarrow(-\infty,\infty]$ be a proper lower semicontinuous $\rho$-weakly convex for some $\rho\in{\mathbb{R}}_{++}$. Proposition C.1 then verifies immediately that, for every $\lambda\in{\mathbb{R}}_{++}$ such that $\lambda\rho<1$, ${\rm s\mbox{-}Prox}_{\lambda f}$ is well-defined everywhere (i.e., a minimizer in (6) exists uniquely for every $x\in{\mathcal{H}}$), and it is $(1-\lambda\rho)$-cocoercive, since the function $\lambda f+(1/2)\left\|\cdot\right\|^{2}$ is strongly convex. This is rather natural from the fact that ${\rm s\mbox{-}Prox}_{f}$ is $(1+\eta)$-cocoercive for an $\eta$-strongly convex function (i.e., the equivalence between statements 1 and 5 in Proposition C.1). Now, the question is whether the converse is true. This is the major point of the following subsection.
If $f-\rho q$ is convex for an arbitrary real number $\rho\in{\mathbb{R}}$, $f$ is $\rho$-convex [vial83]; i.e., an $\eta$-strongly convex function is $\eta$-convex, and a $\eta$-weakly convex function is $(-\eta)$-convex.
E-A Selected Properties and Some Known Results
  • Corollary 11.17 strongly convex implies supercoercivity
  • Proposition 14.15 supercoercivity implies ${\rm dom}~{}f^{*}={\mathcal{H}}$
  • Proposition 16.27 cont f = int dom f = ..
  • Proposition 17.31 continuous –¿ Gateaux diff.
  • Corollary 17.42 Fréchet differentiable
  • Theorem 18.15 (descent lemma)
  • Corollary 18.17 baillon-haddad
  • Example 22.4. strongly convex, strongly monotone
  • Example 22.7. $T$ cocoercive, $T^{-1}$ strongly monotone
  • Proposition 23.13 strongly monotone, cocoercivity of prox
  • Corollary 16.30 $(\partial f)^{-1}=\partial f^{*}$ for $f\in\Gamma_{0}({\mathcal{H}})$
  • Fenchel-Moreau (Theorem 13.37) Let $f$ proper. Then, $f$ is convex and l.s.c. iff $f=f^{**}$
Appendix F General Case
In this section, we consider a general situation when the function $f+\left\|x-\cdot\right\|^{2}/(2\gamma)$ in (6) may have multiple minimizers so that ${\rm s\mbox{-}Prox}_{\gamma}f$ is, in general, a set-valued operator.151515 The proximity operator defined in this way is the same as defined in [46]. In this general case, there is an interplay between maximal cyclic monotonicity and weak convexity as shown below.
F-A Interplay Between Maximal Cyclic Monotonicity and Weak Convexity
Proposition C.2
Given a set-valued operator $T:{\mathcal{H}}\rightarrow 2^{\mathcal{H}}$, the following statements are equivalent.
  1. 1.
    $T$ is maximally cyclically monotone.
  2. 2.
    $T=\partial\psi$ for some convex function $\psi\in\Gamma_{0}({\mathcal{H}})$.
  3. 3.
    $T={\rm s\mbox{-}Prox}_{\phi}$ for some $\phi+(1/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$.
Moreover, if statements 1 – 3 are true, it holds that $\psi^{*}=\phi+(1/2)\left\|\cdot\right\|^{2}$.
Proof.
The equivalence 1) $\Leftrightarrow$ 2) follows by Rockafellar’s cyclic monotonicity theorem [39, Theorem 22.18]. The equivalence 2) $\Leftrightarrow$ 3) can be verified by showing the following equivalence basically:
$~{}T=\partial\psi,~{}\exists\psi\in\Gamma_{0}({\mathcal{H}})$
$\Leftrightarrow$ $~{}T=[\partial(\underbrace{(\psi^{*}-(1/2)\left\|\cdot\right\|^{2})}_{=:\phi}+(1/2)\left\|\cdot\right\|^{2})]^{-1}$
$\hskip 11.00008pt={\rm s\mbox{-}Prox}_{\psi^{*}-(1/2)\left\|\cdot\right\|^{2}},~{}~{}~{}\exists\psi\in\Gamma_{0}({\mathcal{H}}).$
$\Rightarrow$): The first equality is owing to [39, Corollary 16.30], and the second equality can be verified with Lemma 1 as follows: $[\partial((\psi^{*}-(1/2)\left\|\cdot\right\|^{2})+(1/2)\left\|\cdot\right\|^{2})]^{-1}(x)=\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}\Big{[}(\psi^{*}-(1/2)\left\|\cdot\right\|^{2})(y)+\left\|x-y\right\|^{2}/2\Big{]}={\rm s\mbox{-}Prox}_{\psi^{*}-(1/2)\left\|\cdot\right\|^{2}}(x)$ for every $x\in{\mathcal{H}}$. Here, the Fenchel-Moreau theorem [39, Theorem 13.37] ensures that $\psi^{*}\in\Gamma_{0}({\mathcal{H}})(\Leftrightarrow\phi+(1/2)\left\|\cdot\right\|^{2})\in\Gamma_{0}({\mathcal{H}})$.
$\Leftarrow$): Letting $\psi:=(\phi+(1/2)\left\|\cdot\right\|^{2})^{*}$, we have $\psi\in{\Gamma_{0}(\mathcal{H})}$ again by the Fenchel-Moreau theorem [39, Theorem 13.37] so that $T=(\partial(\phi+(1/2)\left\|\cdot\right\|^{2}))^{-1}=\partial[(\phi+(1/2)\left\|\cdot\right\|^{2})^{*}]=\partial\psi$. ∎
F-B Prox Inclusion
Theorem C.1
Let $g:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper function such that $(g+(1/2)\left\|\cdot\right\|^{2})^{*}$ is proper as well. Then, for every $x\in{\mathcal{H}}$, it holds that
${\rm s\mbox{-}Prox}_{g}(x)\subset{\rm s\mbox{-}Prox}_{\tilde{g}}(x),$ (C.7)
where $\tilde{g}:=(g+(1/2)\left\|\cdot\right\|^{2})^{**}-(1/2)\left\|\cdot\right\|^{2}$ so that ${\rm s\mbox{-}Prox}_{\tilde{g}}$ is maximally cyclically monotone.
Proof.
By [39, Proposition 16.10], given a pair $(x,u)\in{\mathcal{H}}^{2}$ of points, we have $x\in[\partial(g+(1/2)\left\|\cdot\right\|^{2})]^{-1}(u)\Leftrightarrow u\in\partial(g+(1/2)\left\|\cdot\right\|^{2})(x)\Rightarrow x\in\partial[(g+(1/2)\left\|\cdot\right\|^{2})^{*}](u)$. Hence, in light of Lemma 1, for every $u\in{\mathcal{H}}$, we have
${\rm s\mbox{-}Prox}_{g}(u)=$ $~{}[\partial(g+(1/2)\left\|\cdot\right\|^{2})]^{-1}(u)$
$\subset$ $~{}\partial[(g+(1/2)\left\|\cdot\right\|^{2})^{*}](u)$
$=$ $~{}[\partial((g+(1/2)\left\|\cdot\right\|^{2})^{**})]^{-1}(u)$
$=$ $~{}[\partial(\tilde{g}+(1/2)\left\|\cdot\right\|^{2})]^{-1}(u)$
$=$ $~{}{\rm s\mbox{-}Prox}_{\tilde{g}}(u).$ (C.8)
Here, the second equality is owing to [39, Corollary 16.30] under $(g+(1/2)\left\|\cdot\right\|^{2})^{*}\in\Gamma_{0}({\mathcal{H}})$ [39, Proposition 13.13]. Since $\tilde{g}+(1/2)\left\|\cdot\right\|^{2}=(g+(1/2)\left\|\cdot\right\|^{2})^{**}\in\Gamma_{0}({\mathcal{H}})$, the maximal cyclic monotonicity of ${\rm s\mbox{-}Prox}_{\tilde{g}}$ follows by Proposition C.2. ∎
By Theorem C.1, if an operator is a selection of a set-valued operator ${\rm s\mbox{-}Prox}_{g}$ for a nonconvex function $g$ that is not even weakly convex, it is also a selection of the set-valued operator ${\rm s\mbox{-}Prox}_{\tilde{g}}$ for the weakly convex function $\tilde{g}:=(g+(1/2)\left\|\cdot\right\|^{2})^{**}-(1/2)\left\|\cdot\right\|^{2}$. This indicates that we can restrict our attention to the class of weakly convex functions in seeking for a better shrinkage operator, for example.
Remark C.1 (On Theorem C.1)
  1. 1.
    The conjugate function $(g+(1/2)\left\|\cdot\right\|^{2})^{*}$ is proper if, for instance, $g$ is bounded below so that $\left\langle{x},{\cdot}\right\rangle-(g+(1/2)\left\|\cdot\right\|^{2})$ is bounded above for every $x\in{\mathcal{H}}$.
  2. 2.
    By the Fenchel-Moreau theorem [39, Theorem 13.37], $\tilde{g}=g$ if and only if $g+(1/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$.
F-C Surjectivity and Weak Convexity
We first prove the following lemma.161616 It is known that convexity of $f:{\mathcal{H}}\rightarrow{\mathbb{R}}$ implies continuity of $f$ when ${\mathcal{H}}$ is finite dimensional [39, Corollary 8.40].
Lemma C.5
Let $g:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper function. Then, the following two statements are equivalent.
  1. 1.
    ${\rm dom}~{}\partial g={\mathcal{H}}$.
  2. 2.
    $g:{\mathcal{H}}\rightarrow{\mathbb{R}}$ is convex and continuous over ${\mathcal{H}}$.
Proof.
1) $\Rightarrow$ 2): By [39, Proposition 16.5] and the Fenchel-Moreau theorem [39, Theorem 13.37], it holds that ${\rm dom}~{}\partial g={\mathcal{H}}\Rightarrow g=g^{**}\Leftrightarrow g\in\Gamma_{0}({\mathcal{H}})$. Hence, invoking [39, Proposition 16.4] and [39, Proposition 16.27], we immediately obtain ${\mathcal{H}}={\rm dom}~{}\partial g\subset{\rm dom}~{}g={\mathcal{H}}\Rightarrow{\rm int}~{}{\rm dom}~{}g={\rm cont}~{}g={\mathcal{H}}$.
2) $\Rightarrow$ 1): Clear from [39, Proposition 16.17(ii)]. ∎
The following result implies that ${\rm s\mbox{-}Prox}_{\varphi}$ in the sense of (6) is surjective if and only if $\varphi$ is continuous and 1-weakly-convex.
Proposition C.3
Assume that $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ can be expressed as $T(x)\in\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}\varphi(y)+\left\|y-x\right\|^{2}/2$ for every $x\in{\mathcal{H}}$ with a proper function $\varphi:{\mathcal{H}}\rightarrow(-\infty,+\infty]$. Consider the following two statements.
  1. 1.
    ${\rm range}~{}T={\mathcal{H}}$.
  2. 2.
    $\varphi+(1/2)\left\|\cdot\right\|^{2}$ is convex and continuous.
Then, 1) $\Rightarrow$ 2). Assume that the function $\varphi(y)+\left\|y-x\right\|^{2}/2$ of $y$ has a unique minimizer so that $T={\rm s\mbox{-}Prox}_{\varphi}$. Then, 1) $\Leftrightarrow$ 2).
Proof.
By the assumption, it holds that ${\mathcal{H}}={\rm range}~{}T\subset{\rm range}~{}[\partial(\varphi+(1/2)\left\|\cdot\right\|^{2})]^{-1}={\rm dom}~{}\partial(\varphi+(1/2)\left\|\cdot\right\|^{2})={\mathcal{H}}$. Hence, by Lemma C.5, ${\rm range}~{}T={\mathcal{H}}$ implies convexity and continuity of $\varphi+(1/2)\left\|\cdot\right\|^{2}$. Under the assumption that $T={\rm s\mbox{-}Prox}_{\varphi}$, by Lemma C.5 again, ${\rm range}~{}T={\rm dom}~{}\partial(\varphi+(1/2)\left\|\cdot\right\|^{2})={\mathcal{H}}$. ∎
F-D $\ell_{0}$ and MC Functions: An Illustrative Example of Theorem C.1
We apply Theorem C.1 to the $\ell_{0}$ norm for ${\mathcal{H}}:={\mathbb{R}}$. Let $g:{\mathbb{R}}\rightarrow{\mathbb{R}}:x\mapsto\begin{cases}0&\mbox{if }x=0\\ 1&\mbox{otherwise.}\end{cases}$ Then, the convex envelope of $g+(\cdot)^{2}/2$ is given by
$(g+(\cdot)^{2}/2)^{**}(x)=\begin{cases}\sqrt{2}\left|x\right|&\mbox{if }\left|x\right|\leq\sqrt{2}\\ x^{2}/2+1&\mbox{if }\left|x\right|>\sqrt{2}\end{cases}$ (C.9)
so that
$\tilde{g}(x)=\begin{cases}\sqrt{2}\left|x\right|-x^{2}/2&\mbox{if }\left|x\right|\leq\sqrt{2}\\ 1&\mbox{if }\left|x\right|>\sqrt{2}.\end{cases}$ (C.10)
The function $\tilde{g}$ is the MC penalty [17] given in (38) for $\lambda_{1}=\lambda_{2}=\sqrt{2}$. It can be shown that
${\rm s\mbox{-}Prox}_{g}:$ $~{}x\mapsto\begin{cases}0&\mbox{if }\left|x\right|<\sqrt{2}\\ \{0,\sqrt{2}\}&\mbox{if }\left|x\right|=\sqrt{2}\\ x&\mbox{if }\left|x\right|>\sqrt{2}\end{cases}$
${\rm s\mbox{-}Prox}_{\tilde{g}}:$ $~{}x\mapsto\begin{cases}0&\mbox{if }\left|x\right|<\sqrt{2}\\ [0,\sqrt{2}]&\mbox{if }\left|x\right|=\sqrt{2}\\ x&\mbox{if }\left|x\right|>\sqrt{2}\end{cases}$ (C.11)
so that ${\rm hard}_{\sqrt{2}}(x)\in{\rm s\mbox{-}Prox}_{g}(x)\subset{\rm s\mbox{-}Prox}_{\tilde{g}}(x)$ for $x\in{\mathbb{R}}$. More precisely, it holds that ${\rm hard}_{\sqrt{2}}=\displaystyle\lim_{\lambda_{2}\downarrow\sqrt{2}}{\rm firm}_{\sqrt{2},\lambda_{2}}$ in the sense of pointwise convergence, and $\tilde{g}=\displaystyle\lim_{\lambda_{2}\downarrow\sqrt{2}}\varphi$ for the $\varphi$ given in (38) with $\lambda_{1}:=\sqrt{2}$ (in the sense of uniform convergence). We remark that, since $\tilde{g}$ is a 1-weakly convex function, Theorem 1 does not apply to $\tilde{g}$, and this justifies the discontinuity of ${\rm s\mbox{-}Prox}_{\tilde{g}}$.
Remark C.2
Suppose that we are given an operator $T:={\rm hard_{\sqrt{2}}}$ and would like to find a weakly convex function $\phi$ such that ${\rm hard_{\sqrt{2}}}(x)\in{\rm s\mbox{-}Prox}_{\phi}(x)$ for every $x\in{\mathbb{R}}$. As monotonicity and cyclic monotonicity are identical in the case of ${\mathcal{H}}:={\mathbb{R}}$, according to Proposition C.2, we wish to obtain the maximal monotone extension of ${\rm hard_{\sqrt{2}}}$, which is actually given by ${\rm s\mbox{-}Prox}_{\tilde{g}}$ in (C.11). We then seek for $\psi^{*}\in\Gamma_{0}({\mathbb{R}})$ such that $\partial\psi^{*}[=(\partial\psi)^{-1}]=({\rm s\mbox{-}Prox}_{\tilde{g}})^{-1}$. In fact, such a function $\psi^{*}$ coincides with $(g+(\cdot)^{2}/2)^{**}$ given in (C.9). Finally, $\phi$ is obtained as $\phi=\psi^{*}-(\cdot)^{2}/2=\tilde{g}$ which is given in (C.10). Thorough discussions about this procedure can be found in [74].
Lemma C.6
Let $g:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper function. Then, given any positive constant $\gamma\in{\mathbb{R}}_{++}$, it holds that
${\rm s\mbox{-}Prox}_{\gamma g}=\left[\partial(g+(\gamma^{-1}/2)\left\|\cdot\right\|^{2})\right]^{-1}\circ(\gamma^{-1}I).$ (C.12)
Proof.
Using Lemma 1, the following equivalence can be verified:
$~{}{\boldsymbol{p}}\in{\rm s\mbox{-}Prox}_{\gamma}g({\boldsymbol{x}})$
$\Leftrightarrow$ $~{}{\boldsymbol{p}}\in\operatornamewithlimits{argmin}_{{\boldsymbol{y}}\in{\mathcal{H}}}\left(\gamma g({\boldsymbol{y}})+\frac{1}{2}\left\|{\boldsymbol{x}}-{\boldsymbol{y}}\right\|^{2}\right)$
$~{}\hskip 20.00003pt=\left[\partial\left(\gamma g+(1/2)\left\|\cdot\right\|^{2}\right)\right]^{-1}({\boldsymbol{x}})$
$\Leftrightarrow$ $~{}{\boldsymbol{x}}\in\partial\left(\gamma g+(1/2)\left\|\cdot\right\|^{2}\right)({\boldsymbol{p}})=\gamma\partial\left(g+(\gamma^{-1}/2)\left\|\cdot\right\|^{2}\right)({\boldsymbol{p}})$
$\Leftrightarrow$ $~{}\gamma^{-1}{\boldsymbol{x}}\in\partial\left(g+(\gamma^{-1}/2)\left\|\cdot\right\|^{2}\right)({\boldsymbol{p}})$
$\Leftrightarrow$ $~{}{\boldsymbol{p}}\in\left[\partial\left(g+(\gamma^{-1}/2)\left\|\cdot\right\|^{2}\right)\right]^{-1}(\gamma^{-1}{\boldsymbol{x}}).$ (C.13)
The following proposition can be verified with Lemma C.6.
Proposition C.4
Let $\phi$ be a 1-weakly convex function, i.e., $\phi+(1/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$. Then, the proximity operator of the $(\tau+1)^{-1}$-weakly convex function $\phi/(\tau+1)$ for $\tau\in{\mathbb{R}}_{++}$ can be expressed as
${\rm s\mbox{-}Prox}_{\phi/(\tau+1)}=\left[{\rm s\mbox{-}Prox}_{\phi}^{-1}+\tau I\right]^{-1}\circ(\tau+1)I.$ (C.14)
Proof.
Owing to the 1-weak convexity of $\phi$, it holds that $\partial(\phi+((\tau+1)/2)\left\|\cdot\right\|^{2})=\partial(\phi+(1/2)\left\|\cdot\right\|^{2})+\partial[(\tau/2)\left\|\cdot\right\|^{2}]=\partial(\phi+(1/2)\left\|\cdot\right\|^{2})+\tau I$. Hence, letting $g:=\phi$ and $\gamma:=1/(\tau+1)\in(0,1)$ in Lemma C.6, we obtain
${\rm s\mbox{-}Prox}_{\phi/(\tau+1)}=$ $~{}\left[\partial(\phi+((\tau+1)/2)\left\|\cdot\right\|^{2})\right]^{-1}\circ(\tau+1)I$
$=$ $~{}\left[\partial(\phi+(1/2)\left\|\cdot\right\|^{2})+\tau I\right]^{-1}\circ(\tau+1)I$
$=$ $~{}\left[{\rm s\mbox{-}Prox}_{\phi}^{-1}+\tau I\right]^{-1}\circ(\tau+1)I.$ (C.15)
Under the setting $\phi:=\tilde{g}$ in Proposition C.4, (C.14) establishes the exact relation between the set-valued operator ${\rm s\mbox{-}Prox}_{\tilde{g}}$ (of which the selection gives the hard shrinkage) and the firm shrinkage operator ${\rm s\mbox{-}Prox}_{\tilde{g}/(\tau+1)}$ for $\lambda_{1}:=\sqrt{2}/(\tau+1)\in(0,\sqrt{2})$ and $\lambda_{2}:=\sqrt{2}$.
F-E Ordered Weighted $\ell_{1}$ Function
We consider the Euclidean case ${\mathcal{H}}:={\mathbb{R}}^{n}$ for $n\geq 2$. Let ${\boldsymbol{w}}\in{\mathbb{R}}_{+}^{n}$ be the weight vector. Given ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, we define $\left|{\boldsymbol{x}}\right|\in{\mathbb{R}}_{+}^{n}$ of which the $i$th component is given by $\left|x_{i}\right|$. Let $\left|{\boldsymbol{x}}\right|_{\downarrow}\in{\mathbb{R}}_{+}^{n}$ denote a sorted version of $\left|{\boldsymbol{x}}\right|$ in the nonincreasing order; i.e., $\big{[}\left|{\boldsymbol{x}}\right|_{\downarrow}\big{]}_{1}\geq\big{[}\left|{\boldsymbol{x}}\right|_{\downarrow}\big{]}_{2}\geq\cdots\geq\big{[}\left|{\boldsymbol{x}}\right|_{\downarrow}\big{]}_{n}$. The ordered weighted $\ell_{1}$ function is defined as
$\Omega_{{\boldsymbol{w}}}({\boldsymbol{x}}):={\boldsymbol{w}}^{\top}\left|{\boldsymbol{x}}\right|_{\downarrow}.$ (C.16)
If the weight is nonincreasing, i.e., $w_{1}\geq w_{2}\geq\cdots\geq w_{n}(\geq 0)$, $\Omega_{{\boldsymbol{w}}}({\boldsymbol{x}})$ is convex, and it is called the ordered weighted $\ell_{1}$ (OWL) norm. For instance, the particular choice of $w_{i}:=\lambda_{1}+\lambda_{2}(n-i)$, $i\in\overline{1,n}$, for some $(\lambda_{1},\lambda_{2})\in{\mathbb{R}}_{++}^{2}$ gives the octagonal shrinkage and clustering algorithm for regression (OSCAR), of which the proximity operator is well defined [58, 60].
If, on the other hand, the weight is not nonincreasing, $\Omega_{{\boldsymbol{w}}}$ is nonconvex, and the proximity operator will be discontinuous. This implies in light of Theorem 1 that $\Omega_{{\boldsymbol{w}}}$ is not even weakly convex. To see this, let us consider the case of $n=2$, and let $(0<)w_{1}<w_{2}$. In this case, the proximity operator is given by
${\rm s\mbox{-}Prox}_{\Omega_{{\boldsymbol{w}}}}({\boldsymbol{x}})=\begin{cases}{\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\uparrow}\right]_{+}&\mbox{if }\left|x_{1}\right|>\left|x_{2}\right|\\ \{{\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\uparrow}\right]_{+},&\\ \hskip 5.0pt{\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\downarrow}\right]_{+}\}&\mbox{if }\left|x_{1}\right|=\left|x_{2}\right|\\ {\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\downarrow}\right]_{+}&\mbox{if }\left|x_{1}\right|<\left|x_{2}\right|\\ \end{cases}$ (C.17)
where ${\boldsymbol{w}}_{\uparrow}:=[w_{1}~{}w_{2}]^{\top}$, ${\boldsymbol{w}}_{\downarrow}:=[w_{2}~{}w_{1}]^{\top}$, and $[\cdot]_{+}$ denotes the componentwise ‘ramp’ function returning $\max\{\cdot,0\}$ for each component of the argument vector. One can see that ${\rm s\mbox{-}Prox}_{\Omega_{{\boldsymbol{w}}}}$ is discontinuous at those points with $\left|x_{1}\right|=\left|x_{2}\right|$.
In analogy with the relation between the $\ell_{0}$ and MC functions discussed in Section F-D, we consider the following modified set-valued operator:
$T({\boldsymbol{x}})=\begin{cases}{\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\uparrow}\right]_{+}&\mbox{if }\left|x_{1}\right|>\left|x_{2}\right|\\ \overline{{\rm conv}}\{{\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\uparrow}\right]_{+},&\\ \hskip 5.0pt{\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\downarrow}\right]_{+}\}&\mbox{if }\left|x_{1}\right|=\left|x_{2}\right|\\ {\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\downarrow}\right]_{+}&\mbox{if }\left|x_{1}\right|<\left|x_{2}\right|,\end{cases}$ (C.18)
which corresponds to ${\rm s\mbox{-}Prox}_{\tilde{g}}$ in (C.11). Here, $\overline{{\rm conv}}\{{\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\uparrow}\right]_{+},{\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\downarrow}\right]_{+}\}=\{\alpha{\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\uparrow}\right]_{+}+(1-\alpha){\rm sgn}({\boldsymbol{x}})\odot\left[\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\downarrow}\right]_{+}\mid\alpha\in[0,1]\}$ is the closed convex hull of the set. It is not difficult to see that $T=\partial\psi$ for some continuous convex function $\psi\in\Gamma_{0}({\mathbb{R}}^{2})$, implying that $T$ is maximally monotone. In the case of $x_{1}>x_{2}\geq w_{2}$, for instance, $\psi({\boldsymbol{x}})=(1/2)\left\|{\boldsymbol{x}}\right\|^{2}-\left\langle{{\boldsymbol{x}}},{{\boldsymbol{w}}_{\uparrow}}\right\rangle$.
Keeping in mind the relation between the set-valued operator ${\rm s\mbox{-}Prox}_{\tilde{g}}$ (related to the hard shrinkage) and the firm shrinkage operator ${\rm s\mbox{-}Prox}_{\tilde{g}/(\tau+1)}$ established under Proposition C.4, we replace ${\rm s\mbox{-}Prox}_{\phi}$ in (C.14) by $T$ to obtain the operator
$\tilde{T}:=\left[T^{-1}+\tau I\right]^{-1}\circ(\tau+1)I.$ (C.19)
One can observe that the operator $T$ is monotone $\Leftrightarrow$ $T^{-1}$ is monotone $\Leftrightarrow$ $T^{-1}+\tau I$ is $\tau$-strongly monotone $\Leftrightarrow$ $(T^{-1}+\tau I)^{-1}$ is $\tau$-cocoercive $\Leftrightarrow$ $\tilde{T}=(T^{-1}+\tau I)^{-1}\circ(\tau+1)I$ is $\tau/(\tau+1)$-cocoercive, or equivalently $(1+1/\tau)$-Lipschitz continuous.
The operator $\tilde{T}$ can be computed as follows:
$\tilde{T}({\boldsymbol{x}})={\rm sign}({\boldsymbol{x}})\odot\begin{cases}(\left|{\boldsymbol{x}}\right|\!-\!{\boldsymbol{w}}_{\uparrow})_{+}&\mbox{if }\left|x_{1}\right|\!-\!\left|x_{2}\right|\!>\!(w_{2}\!-\!w_{1})\tau\\ (\left|{\boldsymbol{x}}\right|\!-\!{\boldsymbol{w}}_{\downarrow})_{+}&\mbox{if}\left|x_{2}\right|\!-\!\left|x_{1}\right|\!>\!(w_{2}\!-\!w_{1})\tau\\ (\left|{\boldsymbol{x}}\right|\!-\!{\boldsymbol{w}}_{\hat{\alpha}({\boldsymbol{x}})})_{+}&\mbox{ otherwise,}\\ \end{cases}$ (C.20)
where ${\boldsymbol{w}}_{\hat{\alpha}({\boldsymbol{x}})}:=\hat{\alpha}({\boldsymbol{x}}){\boldsymbol{w}}_{\uparrow}+(1-\hat{\alpha}({\boldsymbol{x}})){\boldsymbol{w}}_{\downarrow}$ for $\hat{\alpha}({\boldsymbol{x}}):=1/2+(x_{1}-x_{2})/(2(w_{2}-w_{1})\tau)\in(0,1)$. The operator $\tilde{T}$ can be expressed as $\tilde{T}:=\nabla\tilde{\psi}$ for a differentiable convex function $\tilde{\psi}\in\Gamma_{0}({\mathbb{R}}^{2})$. If, for instance, $\left|\left|x_{2}\right|-\left|x_{1}\right|\right|<(w_{2}-w_{1})\tau$ and $\left|{\boldsymbol{x}}\right|-{\boldsymbol{w}}_{\hat{\alpha}({\boldsymbol{x}})}\in{\mathbb{R}}_{++}^{2}$, we have $\tilde{\psi}({\boldsymbol{x}})=(4\tau)^{-1}{\boldsymbol{x}}^{\top}\left[\begin{array}[]{cc}2\tau+1&-1\\ -1&2\tau+1\\ \end{array}\right]{\boldsymbol{x}}-[(w_{1}+w_{2})/2]\left\langle{{\boldsymbol{x}}},{{\boldsymbol{1}}}\right\rangle$. By Theorem 1, $\tilde{T}={\rm s\mbox{-}Prox}_{\tilde{\varphi}}$ for the $1/(\tau+1)$-weakly convex function $\tilde{\varphi}:=\tilde{\psi}^{*}-(1/2)\left\|\cdot\right\|^{2}$.
Appendix G Social Shrinkage
Example C.1 (Social firm-shrinkage)
Consider an $n$-dimensional Euclidean vector ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$. For every index $i\in\mathcal{I}:=\{1,2,\cdots,n\}$, we define the weight matrix ${\boldsymbol{W}}_{i}:={\rm diag}\big{(}w_{1}^{(i)},w_{2}^{(i)},\cdots,w_{n}^{(i)}\big{)}\in{\mathbb{R}}^{n\times n}$, where the nonnegative weights $w_{j}^{(i)}\in[0,1]$, $j\in\mathcal{I}$, satisfy $\sum_{j=1}^{n}w_{j}^{(i)}=1$. Assume here that $w_{i}^{(i)}\neq 0$ for every $i\in\mathcal{I}$. Define the social firm-shrinkage operator
$T^{\rm s\text{-}firm}_{\lambda_{1},\lambda_{2}}:{\mathbb{R}}^{n}\!\rightarrow\!{\mathbb{R}}^{n}:{\boldsymbol{x}}\!\mapsto\!\left[\frac{x_{i}{\rm firm}_{\lambda_{1},\lambda_{2}}\big{(}\big{\|}{\boldsymbol{W}}_{\!\!i}^{1/2}{\boldsymbol{x}}\big{\|}_{2}\big{)}}{\big{\|}{\boldsymbol{W}}_{\!\!i}^{1/2}{\boldsymbol{x}}\big{\|}_{2}}\right]_{i=1}^{n},$ (C.21)
where $\big{\|}{\boldsymbol{W}}_{\!\!i}^{1/2}{\boldsymbol{x}}\big{\|}_{2}=\sqrt{\sum_{j=1}^{n}w_{j}^{(i)}x_{j}^{2}}=\sqrt{\sum_{j\in\mathcal{N}_{i}}w_{j}^{(i)}x_{j}^{2}}$. Here, $\mathcal{N}_{i}:=\{j\in\mathcal{I}\mid w_{j}^{(i)}\neq 0\}$ denotes the neighborhood of the $i$th coefficient $x_{i}$; $j\in\mathcal{N}_{i}(\Leftrightarrow w_{j}^{(i)}\neq 0)$ means that $x_{j}$ belongs to the group of $x_{i}$. Then, the following hold.
  1. 1.
    It holds that
    $T^{\rm s\text{-}firm}_{\lambda_{1},\lambda_{2}}={\boldsymbol{D}}\circ T^{\rm g\text{-}firm}_{\lambda_{1},\lambda_{2}}\circ{\boldsymbol{E}},$ (C.22)
    where ${\boldsymbol{E}}:=[{\boldsymbol{W}}^{1/2}_{1}~{}{\boldsymbol{W}}^{1/2}_{2}~{}\cdots{\boldsymbol{W}}^{1/2}_{n}]^{\top}\in{\mathbb{R}}^{n^{2}\times n}$ and ${\boldsymbol{D}}:=[(w_{1}^{(1)})^{-1/2}{\boldsymbol{e}}_{1}~{}(w_{2}^{(2)})^{-1/2}{\boldsymbol{e}}_{n+2}~{}\cdots~{}(w_{n}^{(n)})^{-1/2}{\boldsymbol{e}}_{n^{2}}]^{\top}\in{\mathbb{R}}^{n\times n^{2}}$ with the standard basis $\{{\boldsymbol{e}}_{i}\}_{i=1}^{n^{2}}$ of ${\mathbb{R}}^{n^{2}}$. Here, $T^{\rm g\text{-}firm}_{\lambda_{1},\lambda_{2}}({\boldsymbol{E}}{\boldsymbol{x}})=(T^{\rm v\text{-}firm}_{\lambda_{1},\lambda_{2}}({\boldsymbol{W}}_{i}^{1/2}{\boldsymbol{x}}))_{i=1}^{n}$.
  2. 2.
    (Non-overlapping case) Let $\mathcal{P}\subset 2^{\mathcal{I}}$ be a partition of the discrete set $\mathcal{I}$,171717 The collection $\{S_{j}\}_{j=1}^{s}$, $s\in\mathcal{I}$, of subsets is a partition of $\mathcal{I}$ if (i) $S_{j}\neq\varnothing$ for every $j\in\overline{1,s}$, (ii) $S_{i}\cap S_{j}=\varnothing$ whenever $i\neq j$, and (iii) $\bigcup_{j\in\overline{1,s}}S_{j}=\mathcal{I}$. and define ${\boldsymbol{w}}_{i}:=[w_{1}^{(i)},w_{2}^{(i)},\cdots,w_{n}^{(i)}]^{\top}$, $i\in\mathcal{I}$. For each set (group) $S\in\mathcal{P}$ from the partition, assume that ${\boldsymbol{w}}_{i}={\boldsymbol{w}}_{j}$ for every pair $(i,j)\in S^{2}$ with ${\rm supp}({\boldsymbol{w}}_{i}):=\{j\in\mathcal{I}\mid w_{j}^{(i)}\neq 0\}=S$. Then, $T^{\rm s\text{-}firm}_{\lambda_{1},\lambda_{2}}$ is the proximity operator of the $(\lambda_{1}/\lambda_{2})$-weakly convex function $\lambda_{1}(\left\|\cdot\right\|_{{\rm g},1})_{\lambda_{2}}\circ{\boldsymbol{E}}$.
Proof.
1) By definition of $T^{\rm g\text{-}firm}_{\lambda_{1},\lambda_{2}}$ and Example 1, we have
$T^{\rm g\text{-}firm}_{\lambda_{1},\lambda_{2}}({\boldsymbol{E}}{\boldsymbol{x}})=\left[\frac{{\boldsymbol{W}}_{\!\!i}^{1/2}{\boldsymbol{x}}~{}{\rm firm}_{\lambda_{1},\lambda_{2}}\big{(}\big{\|}{\boldsymbol{W}}_{\!\!i}^{1/2}{\boldsymbol{x}}\big{\|}_{2}\big{)}}{\big{\|}{\boldsymbol{W}}_{\!\!i}^{1/2}{\boldsymbol{x}}\big{\|}_{2}}\right]_{i=1}^{n}.$ (C.23)
Since $(w_{i}^{(i)})^{-1/2}{\boldsymbol{e}}_{i}^{\top}{\boldsymbol{W}}_{\!\!i}^{1/2}{\boldsymbol{x}}=x_{i}$, (C.22) is verified.
2) We let $\Omega:=(\left\|\cdot\right\|_{{\rm g},1})_{\lambda_{2}}$. Keeping in mind $T^{\rm g\text{-}firm}_{\lambda_{1},\lambda_{2}}={\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}$ (Example 2), we prove the claim by the following steps.
  1. (a)
    ${\rm s\mbox{-}Prox}_{\lambda_{1}\Omega\circ{\boldsymbol{E}}}={\boldsymbol{E}}^{\top}\circ{\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}\circ{\boldsymbol{E}}.$
  2. (b)
    ${\boldsymbol{D}}\circ{\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}\circ{\boldsymbol{E}}={\boldsymbol{E}}^{\top}\circ{\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}\circ{\boldsymbol{E}}$.
(a) We use the following lemma.
Lemma C.7
Let ${\mathcal{H}}$ and $\mathcal{K}$ be real Hilbert spaces equipped with the norms $\left\|\cdot\right\|_{{\mathcal{H}}}$ and $\left\|\cdot\right\|_{\mathcal{K}}$, respectively. Also, let $f:\mathcal{K}\rightarrow(-\infty,+\infty]$ be a function such that $f\circ L+((1-\beta)/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ for some $\beta\in(0,1)$, where $L:{\mathcal{H}}\rightarrow\mathcal{K}$ be a bounded linear operator such that $L^{*}L={\rm Id}$. Assume that $f+\left\|Lx-\cdot\right\|^{2}/2$ has a (unique) minimizer on ${\rm range}~{}L$. Then, it holds that ${\rm s\mbox{-}Prox}_{f\circ L}=L^{*}\circ{\rm s\mbox{-}Prox}_{f}\circ L$.
By ${\boldsymbol{E}}^{\top}{\boldsymbol{E}}=\sum_{i=1}^{n}{\boldsymbol{E}}_{i}^{\top}{\boldsymbol{E}}_{i}={\boldsymbol{I}}_{n}$, in light of Lemma C.7, it is sufficient to show the following inclusion:
${\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}({\boldsymbol{E}}{\boldsymbol{x}})\in{\rm range}~{}{\boldsymbol{E}},~{}\forall{\boldsymbol{x}}\in{\mathbb{R}}^{n}.$ (C.24)
Fix an ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ arbitrarily. Owing to the group-wise separability of $\left\|\cdot\right\|_{{\rm g},1}$, it can be verified that
$J({\boldsymbol{y}}):=\Omega({\boldsymbol{y}})+\frac{1}{2\lambda_{1}}\left\|{\boldsymbol{E}}{\boldsymbol{x}}-{\boldsymbol{y}}\right\|_{2}^{2}$
$=\sum_{i=1}^{n}\Big{(}\underbrace{(\left\|\cdot\right\|_{2})_{\lambda_{2}}({\boldsymbol{y}}_{i})+\frac{1}{2\lambda_{1}}\big{\|}{\boldsymbol{W}}_{i}^{1/2}{\boldsymbol{x}}-{\boldsymbol{y}}_{i}\big{\|}_{2}^{2}}_{=:J_{i}({\boldsymbol{y}}_{i})}\Big{)},$ (C.25)
where ${\boldsymbol{y}}_{i}\in{\mathbb{R}}^{n}$ such that ${\boldsymbol{y}}=[{\boldsymbol{y}}_{1}^{\top}~{}{\boldsymbol{y}}_{2}^{\top}~{}\cdots~{}{\boldsymbol{y}}_{n}^{\top}]^{\top}$. The minimization of the function in (C.25) can be considered for each ${\boldsymbol{y}}_{i}$ separately. Because of the $(\lambda_{1}/\lambda_{2})$-weak convexity of $\lambda_{1}(\left\|\cdot\right\|_{2})_{\lambda_{2}}$, each function $J_{i}({\boldsymbol{y}}_{i}):=(\left\|\cdot\right\|_{2})_{\lambda_{2}}({\boldsymbol{y}}_{i})+\big{\|}{\boldsymbol{W}}_{i}^{1/2}{\boldsymbol{x}}-{\boldsymbol{y}}_{i}\big{\|}_{2}^{2}/(2\lambda_{1})$ has a unique minimizer, say $\hat{{\boldsymbol{y}}}_{i}$, which lies in ${\rm span}\{{\boldsymbol{W}}_{i}^{1/2}{\boldsymbol{x}}\}$; this can be verified by the same argument as in the proof of Example 1. Moreover, for every pair $(i,j)\in S^{2}$ in each group $S\in\mathcal{P}$, the associated functions $J_{i}$ and $J_{j}$ coincide with each other (as ${\boldsymbol{W}}_{i}={\boldsymbol{W}}_{j}$ by the assumption). Hence, their minimizers coincide as well, i.e., $\hat{{\boldsymbol{y}}}_{i}=\hat{{\boldsymbol{y}}}_{j}\in{\rm span}\{{\boldsymbol{W}}_{i}^{1/2}{\boldsymbol{x}}\}$, meaning that $[\hat{{\boldsymbol{y}}}_{i}^{\top}~{}\hat{{\boldsymbol{y}}}_{j}^{\top}]^{\top}=[{\boldsymbol{W}}_{i}^{1/2}~{}{\boldsymbol{W}}_{j}^{1/2}]^{\top}(\xi{\boldsymbol{x}})\in{\rm range}~{}[{\boldsymbol{W}}_{i}^{1/2}~{}{\boldsymbol{W}}_{j}^{1/2}]^{\top}$ for some $\xi\in{\mathbb{R}}$. Based on this, it can be verified that ${\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}({\boldsymbol{E}}{\boldsymbol{x}})=\operatornamewithlimits{argmin}_{{\boldsymbol{y}}\in{\mathbb{R}}^{n}}J({\boldsymbol{y}})\in{\rm range}~{}{\boldsymbol{E}}$.
(b) Owing to ${\boldsymbol{E}}^{\top}{\boldsymbol{E}}={\boldsymbol{I}}_{n}$, ${\boldsymbol{E}}{\boldsymbol{E}}^{\top}$ gives the projection matrix onto ${\rm range}~{}{\boldsymbol{E}}$, and thus the inclusion in (C.24) implies that ${\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}({\boldsymbol{E}}{\boldsymbol{x}})={\boldsymbol{E}}{\boldsymbol{E}}^{\top}\big{(}{\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}({\boldsymbol{E}}{\boldsymbol{x}})\big{)}$, from which together with ${\boldsymbol{D}}{\boldsymbol{E}}={\boldsymbol{I}}_{n}$ it follows that ${\boldsymbol{D}}\circ{\rm s\mbox{-}Prox}_{\lambda_{1}(\left\|\cdot\right\|_{{\rm g},1})_{\lambda_{2}}}\circ{\boldsymbol{E}}={\boldsymbol{D}}{\boldsymbol{E}}{\boldsymbol{E}}^{\top}\circ{\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}\circ{\boldsymbol{E}}={\boldsymbol{E}}^{\top}\circ{\rm s\mbox{-}Prox}_{\lambda_{1}\Omega}\circ{\boldsymbol{E}}$. ∎
Proof of Lemma C.7: The claim can be verified as follows:
${\rm s\mbox{-}Prox}_{f\circ L}(x)=$ $~{}\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}\Big{(}f(Ly)+\frac{1}{2}\left\|x-y\right\|_{{\mathcal{H}}}^{2}\Big{)}$
$=$ $~{}\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}\Big{(}f(Ly)+\frac{1}{2}\left\|Lx-Ly\right\|_{\mathcal{K}}^{2}\Big{)}$
$=$ $~{}L^{*}\operatornamewithlimits{argmin}_{z\in{\rm range}~{}L}\Big{(}f(z)+\frac{1}{2}\left\|Lx-z\right\|_{\mathcal{K}}^{2}\Big{)}$
$=$ $~{}L^{*}\operatornamewithlimits{argmin}_{z\in\mathcal{K}}\Big{(}f(z)+\frac{1}{2}\left\|Lx-z\right\|_{\mathcal{K}}^{2}\Big{)}$
$=$ $~{}L^{*}{\rm s\mbox{-}Prox}_{f}(Lx),$ (C.26)
where the fourth equality is owing to the assumption. $\blacksquare$
Example C.1 is a firm-shrinkage version of the windowed group lasso [kowalski13]. In fact, the windowed group lasso can be regarded as an asymptotic case of $T^{\rm s\text{-}firm}_{\lambda_{1},\lambda_{2}}$ with $\lambda_{2}\rightarrow+\infty$ (which makes $\beta\downarrow 0$).
G-A Averagedness Under Similar Transformation
The expression in (C.22) raises the question of when an operator $T$ with linear compositions from both sides can be cocoercive, or, more directly, be the proximity operator of a weakly convex function. This question is addressed below.
Lemma C.8
Let $X$ and ${\mathcal{H}}$ be finite-dimensional real Hilbert spaces and $A:X\rightarrow{\mathcal{H}}$ a bounded linear operator such that $A^{\dagger}A={\rm Id}_{X}$. Here, $A^{\dagger}$ is the generalized (Moore-Penrose) inverse of $A$, and ${\rm Id}_{X}$ is the identity operator in $X$. Define $\kappa:=\sigma_{\max}(A)/\sigma_{\min}(A)\in[1,+\infty)$, where $\sigma_{\max}(\cdot)$ and $\sigma_{\min}(\cdot)$ denote the maximal and minimal singular values, respectively. Given a (nonlinear) operator $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$, we define $\tilde{T}:=A^{\dagger}\circ T\circ A:X\rightarrow X$. Then, the following hold.
  1. 1.
    Assume that $T$ is $\alpha$-averaged nonexpansive for $\alpha\in(0,1)$. Then, $\chi^{-1}\tilde{T}$ is $\tilde{\alpha}$-averaged nonexpansive, where $\chi:=\alpha(\kappa-1)+1\in[1,+\infty)$ and $\tilde{\alpha}:=\alpha\kappa/\chi\in(0,1)$. In particular, if $\kappa=1$ (i.e., all the singular values of $A$ coincide), $\tilde{T}$ is $\alpha$-averaged nonexpansive.
  2. 2.
    Assume that $T$ is $\gamma$-cocoercive for $\gamma\in{\mathbb{R}}_{++}$. Then, $[2\gamma/(\kappa+1)]\tilde{T}$ is $\tilde{\alpha}$-averaged nonexpansive, where $\tilde{\alpha}:=\kappa/(\kappa+1)\in[1/2,1)$. In particular, if $\kappa=1$, $\tilde{T}$ is $\gamma$-cocoercive.
Proof.
1): By the assumption, there exists a nonexpansive operator $N:{\mathcal{H}}\rightarrow{\mathcal{H}}$ such that $T=(1-\alpha){\rm Id}+\alpha N$. Hence, we have
$\chi^{-1}\tilde{T}=$ $~{}\chi^{-1}A^{\dagger}\circ[(1-\alpha){\rm Id}+\alpha N]\circ A$
$=$ $~{}\chi^{-1}(1-\alpha)A^{\dagger}A+\chi^{-1}\alpha A^{\dagger}\circ N\circ A$
$=$ $~{}(1-\tilde{\alpha}){\rm Id}+\tilde{\alpha}\tilde{N},$ (C.27)
where $\tilde{N}:=\kappa^{-1}A^{\dagger}\circ N\circ A$ is nonexpansive as verified by
$\left\|\tilde{N}(x)-\tilde{N}(y)\right\|=$ $~{}\kappa^{-1}\left\|A^{\dagger}(N\circ A(x)-N\circ A(y))\right\|$
$\leq$ $~{}\kappa^{-1}\sigma_{\min}^{-1}(A)\left\|N(Ax)-N(Ay)\right\|$
$\leq$ $~{}\sigma_{\max}^{-1}(A)\left\|A(x-y)\right\|$
$\leq$ $~{}\left\|x-y\right\|.$ (C.28)
Here, the nonexpansiveness of $N$ is used to verify the second inequality. When $\kappa=1$, we have $\chi=1$ and $\tilde{\alpha}=\alpha$, which verifies the assertion.
2): Applying Lemma C.8.1 to the $\alpha:=1/2$-averaged nonexpansive operator $T_{1}:=\gamma T$, $\chi_{1}^{-1}A^{\dagger}\circ T_{1}\circ A$ is $\tilde{\alpha}_{1}$-averaged nonexpansive with $\chi_{1}=(\kappa+1)/2$ and $\tilde{\alpha}=\kappa/(2\chi_{1})=\kappa/(\kappa+1)$. The assertion is verified by observing that $\chi_{1}^{-1}A^{\dagger}\circ T_{1}\circ A=[2\gamma/(\kappa+1)]A^{\dagger}\circ T\circ A$. ∎
Lemma B.1 is reproduced by Lemma C.8.1. Specifically, Lemma B.1.1 is reproduced by letting $A:=c{\rm Id}$, while Lemma B.1.2 is reproduced by letting $A:=\beta{\rm Id}$ and inserting $\beta T$ into the $T$ of Lemma C.8.1.
Proposition C.5
Let $X$ and ${\mathcal{H}}$ be finite-dimensional real Hilbert spaces and $A:X\rightarrow{\mathcal{H}}$ a bounded linear operator such that $A^{*}A={\rm Id}_{X}$, where $A^{*}$ is the adjoint operator of $A$. Let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ be $\alpha$-averaged nonexpansive (resp. $\gamma$-cocoercive), where $\alpha\in(0,1)$ and $\gamma\in{\mathbb{R}}_{++}$. Then, $\tilde{T}:=A^{*}\circ T\circ A$ is $\alpha$-averaged nonexpansive (resp. $\gamma$-cocoercive). If, moreover, $T=\nabla\psi$ for some $\psi\in\Gamma_{0}({\mathcal{H}})$, then $\tilde{T}=\nabla(\psi\circ A)$.
Proof.
By $A^{*}A={\rm Id}_{X}$, it holds that (i) $A^{\dagger}=(A^{*}A)^{-1}A^{*}=A^{*}$ [39, Example 3.29], and (ii) $\kappa=1$ (because $\sigma_{1}=\sigma_{2}=\cdots=\sigma_{n}=1$). Hence, the assertions are verified with Lemma C.8. ∎
Example C.2
For ${\mathcal{H}}:={\mathbb{R}}^{n}$, let $A:={\boldsymbol{E}}:=[{\boldsymbol{E}}_{1}~{}{\boldsymbol{E}}_{2}~{}\cdots{\boldsymbol{E}}_{n}]^{\top}\in{\mathbb{R}}^{n^{2}\times n}$, where ${\boldsymbol{E}}_{i}:=\Big{(}{\rm diag}\big{(}w_{1}^{(i)},w_{2}^{(i)},\cdots,w_{n}^{(i)}\big{)}\Big{)}^{1/2}\in{\mathbb{R}}^{n\times n}$, $i\in\overline{1,n}:=\{1,2,\cdots,n\}$, with nonnegative weights $w_{j}^{(i)}\in[0,1]$, $j\in\overline{1,n}$, such that $\sum_{j=1}^{n}w_{j}^{(i)}=1$. In this case, $A^{*}A={\boldsymbol{E}}^{\top}{\boldsymbol{E}}=\sum_{i=1}^{n}{\boldsymbol{E}}_{i}^{\top}{\boldsymbol{E}}_{i}={\boldsymbol{I}}_{n}$. Also let $T:={\rm s\mbox{-}Prox}_{\Omega}$ for a convex function $\Omega\in\Gamma_{0}({\mathbb{R}}^{n^{2}})$. Hence, by Proposition C.5, $\tilde{T}={\boldsymbol{E}}^{\top}\circ{\rm s\mbox{-}Prox}_{\Omega}\circ{\boldsymbol{E}}$ is $(1/2)$-averaged nonexpansive, and $\tilde{T}=\nabla\hskip 1.99997pt(^{1}(\Omega^{*})\circ{\boldsymbol{E}})$ since ${\rm s\mbox{-}Prox}_{\Omega}=\nabla\hskip 1.99997pt^{1}(\Omega^{*})$. Invoking Theorem 1, we have $\tilde{T}={\rm s\mbox{-}Prox}_{\varphi}$ with $\varphi:=(^{1}(\Omega^{*})\circ{\boldsymbol{E}})^{*}-(1/2)\left\|\cdot\right\|^{2}=[{\boldsymbol{E}}^{\top}\triangleright(\Omega+(1/2)\left\|\cdot\right\|^{2})]\hskip 1.99997pt{\breve{}}-(1/2)\left\|\cdot\right\|^{2}$, where $\triangleright$ denotes the infimal postcomposition and  ${\breve{}}$ denotes the lower semicontinuous convex envelope.181818 Given a function $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ and an operator $L:{\mathcal{H}}\rightarrow X$ from ${\mathcal{H}}$ to a vector space $X$, the infimal postcomposition of $f$ by $L$ is defined by $L\triangleright f:X\rightarrow(-\infty,+\infty]:y\mapsto\displaystyle\inf_{x\in{\mathcal{H}},~{}Lx=y}f(x)$. The lower semicontinuous convex envelope $\breve{f}$ of $f$ is the supremum of lower semicontinuous convex minorants of $f$. Let, for instance, $\Omega:=\left\|\cdot\right\|_{2,1}:{\mathbb{R}}^{n^{2}}\ni[{\boldsymbol{x}}_{1}^{\top}~{}{\boldsymbol{x}}_{2}^{\top}~{}\cdots~{}{\boldsymbol{x}}_{n}^{\top}]^{\top}\mapsto\sum_{i=1}^{n}\left\|{\boldsymbol{x}}_{i}\right\|_{2}$, where ${\boldsymbol{x}}_{i}\in{\mathbb{R}}^{n}$, $i\in\overline{1,n}$. In this case, $\tilde{T}$ is the so-called orth-WG-lasso, which is a particular instance of the social shrinkage. Here, “WG” stands for “windowed group”.
Remark C.3
Lemma C.8 holds if $A^{\dagger}$ is replaced by a bounded linear operator $B:{\mathcal{H}}\rightarrow X$ such that $BA={\rm Id}_{X}$ with $\kappa:=\sigma_{\max}(A)\sigma_{\max}(B)\in[1,+\infty)$.191919 Note that $\sigma_{\max}(A)\sigma_{\max}(B)\geq\sigma_{\max}(AB)$ because $\left\|A\right\|_{*}\left\|B\right\|_{*}\geq\left\|AB\right\|_{*}$ for every operator norm $\left\|\cdot\right\|_{*}$ (submultiplicativity of operator norms). For the case of Example C.2, we define ${\boldsymbol{D}}:=[(w_{1}^{(1)})^{-1/2}{\boldsymbol{e}}_{1}~{}(w_{2}^{(2)})^{-1/2}{\boldsymbol{e}}_{n+2}~{}\cdots~{}(w_{n}^{(n)})^{-1/2}{\boldsymbol{e}}_{n^{2}}]^{\top}\in{\mathbb{R}}^{n\times n^{2}}$, where $\{{\boldsymbol{e}}_{i}\}_{i=1}^{n^{2}}\subset{\mathbb{R}}^{n^{2}}$ is the standard basis of ${\mathbb{R}}^{n^{2}}$. It can be verified that ${\boldsymbol{D}}{\boldsymbol{E}}={\boldsymbol{I}}_{n}$. The operator $\tilde{T}_{{\boldsymbol{D}}}={\boldsymbol{D}}\circ{\rm s\mbox{-}Prox}_{\Omega}\circ{\boldsymbol{E}}$ produces the social shrinkage operators of WG-lasso, windowed elitist (WE) lasso, and persistent elitist (PE) lasso, depending on the choice of $\Omega$ [kowalski13, Proposition 3]. In this case, $\kappa=\sigma_{\max}({\boldsymbol{D}})\sigma_{\max}({\boldsymbol{E}})=\max\{(w_{i}^{(i)})^{-1/2}\}_{i=1}^{n}\in[1,+\infty)$.
We assume that $\kappa>1$, which is satisfied excluding the trivial case of $w_{1}^{(1)}=w_{2}^{(2)}=\cdots=w_{n}^{(n)}=1$. Since ${\rm s\mbox{-}Prox}_{\Omega}$ is (1/2)-averaged nonexpansive (i.e., $\alpha=1/2$), $\beta=(\kappa+1)/2\in(1,+\infty)$ and $\tilde{\alpha}=\kappa/(\kappa+1)\in(1/2,1)$, which means that cocoercivity of $\tilde{T}_{{\boldsymbol{D}}}$ is not guaranteed. The case of “non-overlapping” groups is exceptional here. Specifically, let $\mathcal{P}\subset 2^{\mathcal{I}}$ be a partition of the discrete set $\mathcal{I}$,202020 The collection $\{S_{j}\}_{j=1}^{s}$, $s\in\mathcal{I}$, of subsets is a partition of $\mathcal{I}$ if (i) $S_{j}\neq\varnothing$ for every $j\in\overline{1,s}$, (ii) $S_{i}\cap S_{j}=\varnothing$ whenever $i\neq j$, and (iii) $\bigcup_{j\in\overline{1,s}}S_{j}=\mathcal{I}$. and define ${\boldsymbol{w}}_{i}:=[w_{1}^{(i)},w_{2}^{(i)},\cdots,w_{n}^{(i)}]^{\top}$, $i\in\mathcal{I}$. For each set (group) $S\in\mathcal{P}$ from the partition, we let ${\boldsymbol{w}}_{i}={\boldsymbol{w}}_{j}$ for every pair $(i,j)\in S^{2}$ with ${\rm supp}({\boldsymbol{w}}_{i}):=\{j\in\mathcal{I}\mid w_{j}^{(i)}\neq 0\}=S$. In this non-overlapping case, it can be verified that ${\rm s\mbox{-}Prox}_{\Omega}({\boldsymbol{E}}{\boldsymbol{x}})\in{\rm range}~{}{\boldsymbol{E}}$. Hence, because, given ${\boldsymbol{u}}\in{\mathbb{R}}^{n^{2}}$, ${\boldsymbol{u}}\in{\rm range}~{}{\boldsymbol{E}}(\Leftrightarrow{\boldsymbol{u}}={\boldsymbol{E}}{\boldsymbol{E}}^{\top}{\boldsymbol{u}})\Rightarrow{\boldsymbol{D}}{\boldsymbol{u}}={\boldsymbol{E}}^{\top}{\boldsymbol{u}}$ [kowalski13], it follows that ${\boldsymbol{D}}\circ{\rm s\mbox{-}Prox}_{\Omega}\circ{\boldsymbol{E}}={\boldsymbol{E}}^{\top}\circ{\rm s\mbox{-}Prox}_{\Omega}\circ{\boldsymbol{E}}$ and thus $\tilde{T}_{{\boldsymbol{D}}}={\rm s\mbox{-}Prox}_{\varphi}$ for the $\varphi$ given in Example C.2. In the overlapping case, on the other hand, the WG-lasso shrinkage operator $\tilde{T}_{{\boldsymbol{D}}}$ for the $\ell_{2,1}$ norm $\Omega:=\left\|\cdot\right\|_{2,1}$ cannot be expressed as ${\rm s\mbox{-}Prox}_{\varphi}$ for every function $\varphi$ [45].